### v1

Abstract
Real-world contracts are often ambiguous. Recent work by Dütting et al. (EC 2023, Econometrica 2024) models ambiguous contracts as a collection of classic contracts, with the agent choosing an action that maximizes his worst-case utility. In this model, optimal ambiguous contracts have been shown to be “simple” in that they consist of single-outcome payment (SOP) contracts, and can be computed in polynomial-time. However, this simplicity is challenged by the potential need for many classic contracts. Motivated by this, we explore succinct ambiguous contracts, where the ambiguous contract is restricted to consist of at most $k$ classic contracts. Unlike in the unrestricted case, succinct ambiguous contracts are no longer composed solely of SOP contracts, making both their structure and computation more complex.
We show that, despite this added complexity, optimal succinct ambiguous contracts are governed by a simple divide-and-conquer principle, showing that they consist of “shifted min-pay contracts” for a suitable partition of the actions. This structural insight implies a characterization of implementability by succinct ambiguous contracts, and can be leveraged to devise an algorithm for the optimal succinct ambiguous contract. While this algorithm is polynomial for $k$ sufficiently close to $n$, for smaller values of $k$, this algorithm is exponential, and we show that this is inevitable (unless P=NP) by establishing NP-hardness for any constant $k$, or $k=\beta n$ for some $\beta\in(0,1)$. Finally, we introduce the succinctness gap measure to quantify the loss incurred due to succinctness, and provide upper and lower bounds on this gap. Interestingly, in the case where we are missing just a single contract from the number sufficient to obtain the utility of the unrestricted case, the principal’s utility drops by a factor of $2$, and this is tight.
1 Introduction
Contract design is a cornerstone of microeconomic theory Holmström (1979); Ross (1973); Mirrlees (1975); Grossman and Hart (1983), as evidenced by the 2016 Nobel Prize in Economics awarded to Hart and Holmström (The Royal Swedish Academy of Sciences, 2016). Just as mechanism design is fundamental to markets for goods, contract design serves a crucial role in markets for services. In recent years, it has also gained growing interest within the computer science community, for a recent survey see Dütting et al. (2024a).
Real-world contracts are frequently ambiguous, as reflected in the inherently ambiguous terms they often contain. A software development contract might stipulate that the code be written in a “maintainable and efficient” manner, an employment contract might state that an employee must demonstrate “professionalism and commitment,” and a university tenure policy may require a faculty member to show “impactful contributions in their field.” In each of these cases, the precise meaning of the terms is open to interpretations.
The potential role of ambiguity in contracts was first noted by Bernheim and Whinston (1998). More recently, Dütting et al. (2024) proposed a formal model of ambiguous contracts, inspired by a similar framework for ambiguous auctions (Di Tillio et al., 2017).
Their model builds on a standard principal-agent framework with $n$ actions and $m$ outcomes. Each action $i$ has an associated cost $c_{i}$, and an outcome distribution represented by a vector $p_{i}$, where $p_{ij}$ denotes the probability of outcome $j$ under action $i$. The principal observes only the realized outcomes, and, to incentivize the agent, designs a contract — a payment function $t:[m]\rightarrow\mathbb{R}_{+}$ that specifies a payment for each possible outcome. The agent’s utility for action $i$ is given by the expected payment for action $i$ minus the action’s cost $c_{i}$.
The key innovation in their model is the introduction of ambiguity. An ambiguous contract $\tau=\{t^{1},t^{2},\ldots\}$ consists of a collection of payment functions, with the agent evaluating each action based on the worst-case utility over the contract’s support. This max-min evaluation captures the agent’s ambiguity aversion (Schmeidler, 1989; Gilboa and Schmeidler, 1993). To ensure the credibility of the principal’s offer, an ambiguous contract must satisfy a consistency condition, which requires that the principal be indifferent among all contracts within the ambiguous contract, given the agent’s best response. By leveraging this form of ambiguity, the principal can typically incentivize actions at a lower payment, or implement actions that would otherwise be infeasible to implement with a classic contract.
Dütting et al. (2024) have shown that optimal ambiguous contracts exhibit structural simplicity, as they are composed of simple single-outcome payment (SOP) functions — payment functions that assign payment to only a single outcome. However, while each SOP contract is inherently simple, a contract composed of many such functions can become increasingly complex. Under the MLRP regularity condition, two classic contracts are sufficient to fully exploit the benefits of ambiguity. In contrast, in the general case, achieving the full potential of ambiguous contracts may require (roughly) as many contracts as there are actions or outcomes.
Succinct Ambiguous Contracts.
A natural way to manage this growing complexity is to focus on ambiguous contracts with small support. Specifically, a $k$-ambiguous contract is an ambiguous contract consisting of at most $k$-classic contracts. Indeed, While it may be unrealistic for a contract to offer a vast number of different compensation schemes, it is quite common for contracts to specify a limited number of possibilities.
For example, sales commission contracts often define multiple bonus structures: one for meeting a baseline sales target and another, higher bonus, for exceeding a stretch goal. However, the exact criteria for receiving the higher bonus may not be fully transparent, as it may depend on subjective performance evaluations or discretionary factors determined by management.
It is not difficult to see that when ambiguous contracts are restricted to a limited size, the optimal contract is no longer composed solely of SOP classic contracts. Since the SOP structure was the key property enabling feasible computation of the optimal ambiguous contract, this leads to two natural questions:
  • What is the structure of an optimal $k$-ambiguous contract?
  • What is the computational complexity of computing the optimal $k$-ambiguous contract?
Beyond algorithmic and computational concerns, a third question focuses on quantifying the principal’s loss due to the restriction to succinct contracts:
  • How much utility does the principal lose when the number of contracts is limited to $k$, compared to an unrestricted ambiguous contract?
In this work, we address all three questions.
1.1 Our Results
Structure and Algorithms.
Our first result demonstrates that, optimal succinct ambiguous contracts follow a straightforward divide-and-conquer principle, revealing that they are composed of “shifted min-pay contracts” over an appropriately chosen partition of the actions. A shifted min-pay contract is obtained by solving the familiar MIN-PAY-LP$(i)$ (see Figure 1) for a target action $i$ and a subset of the actions, and by subsequently increasing the resulting payments by a fixed amount.
Theorem (Theorem 3.1). For any $k$, there is an optimal IC $k$-ambiguous contract $\langle\tau,i^{\star}\rangle$ with $\tau=\{t^{1},\ldots,t^{k}\}$ that takes the following form:
  • There is a partition of the actions $[n]\setminus\{i^{\star}\}$ into $k$ sets $\{S^{1},\ldots,S^{k}\}$.
  • For each $\ell\in[k]$, contract $t^{\ell}$ is a shifted min-pay contract for action $i^{\star}$ for the subinstance $\mathcal{I}_{\{i^{\star}\}\cup S^{\ell}}$, obtained by restricting the original instance to actions $\{i^{\star}\}\cup S^{\ell}$.
To prove this characterization, we first define the concept of a set of $k$ payment functions that “protect” a target action against a partition of the remaining actions into $k$ sets. We then show how such a protective set of payment functions can be turned into an IC $k$-ambiguous contract. Using these observations, we show how to compute an optimal IC $k$-ambiguous contract for a given target action and a fixed partition of the remaining actions into $k$ sets. We conclude that, by iterating over all actions and all partitions of the remaining actions into $k$ sets, we obtain an optimal IC $k$-ambiguous contract.
Beyond the algorithm derived from this structural insight, it also provides a characterization of the implementability of an action $i$ by an IC $k$-ambiguous contract, meaning that there exists a $k$-ambiguous contract under which the agent prefers action $i$ over any other action.
Proposition (Proposition 3.1). An action $i$ in an instance $\mathcal{I}$ is implementable by a $k$-ambiguous contract if and only if there exists a partition $\mathcal{S}$ of $[n]\setminus\{i\}$ to $k$ sets, such that for every set $S\in\mathcal{S}$, action $i$ is implementable by a classic contract in subinstance $\mathcal{I}_{\{i\}\cup S}$.
This characterization suggests that increasing $k$ from $1$ to $n-1$ gradually relaxes the implementability constraints, with classic contracts and unrestricted ambiguous contracts serving as the two endpoints of this spectrum.
Notably, both our implementability and algorithmic results generalize the characterizations established in the unrestricted setting of Dütting et al. (2024). Moreover, our arguments offer alternative proofs for these characterizations in the unrestricted case.
Computational Complexity.
Our algorithm for finding an optimal $k$-ambiguous contract iterates over all actions $i\in[n]$. For each action $i$, it further iterates over all partitions of $[n]\setminus\{i\}$ into $k$ sets. Thus, its running time is proportional to the number of partitions of a set of size $n-1$ into $k$ non-empty sets, which is given by the Stirling number of the second kind {${}{n-1}{k}$. In the two extremes, for $k=n-2$, {${}{n-1}{n-2}=O(n^{2})$, while for $k=2$, {${}{n-1}{2}=O(2^{n})$. More generally, this algorithm is super-polynomial-in-$n$, unless $k=n-O(1)$.
This raises the natural question of whether a more efficient algorithm could bypass this complexity. We establish the hardness of the optimal $k$-ambiguous contract using a reduction from the Makespan Minimization problem, showing that the optimal $k$-ambiguous contract problem is NP-hard for any $k$ for which finding a partition of $n$ numbers into $k$ sets that minimizes the makespan is NP-hard. This includes any constant $k$ or any $k=\beta\cdot n$ for $\beta\in(0,1)$.
Theorem (Theorem 4.1). The $k(n)$-ambiguous contract problem is NP-hard for all functions $k:[n]\rightarrow[n]$ such that $(n,k(n))$-Makespan Minimization is NP-hard.
Besides showing hardness, our reduction reveals a interesting structural parallel between the two problems. It shows how any instance of the Makespan Minimization problem can be turned into an instance of the contracting problem, such that an optimal $k$-ambiguous contract yields an optimal solution to the Makespan Minimization instance.
Succinctness Gap.
We introduce the notion of a succinctness gap, which measures the principal’s loss in utility when limited to $k$-ambiguous contracts. Specifically, the succinctness gap of an instance is defined as the ratio between the principal’s optimal utility under $k$-ambiguous contracts and their optimal utility under unconstrained ambiguous contracts. The succinctness gap of a class of instances is the supremum succinctness gap across all instances in the class.
Prior work has shown that considering ambiguous contracts of size at most $n-1$ is without loss (Dütting et al., 2024). At the other extreme, an ambiguous contract of size $1$ is a classic contract, for which the loss in principal’s utility compared to the optimal unconstrained ambiguous contract was shown to be tightly $n-1$ (Dütting et al., 2024). We establish upper and lower bounds for this gap for all values of $k$ within this spectrum. Our first result shows a general lower bound on the succinctness gap, for all values of $k$. It shows that the principal’s utility from a $k$-ambiguous contract is at least a $\frac{1}{n-k}$ fraction of the principal’s utility in the unrestricted case.
Theorem (Theorem 5.1). For any instance with $n$ actions, for any $k=1,\ldots,n-1$, there exists an IC $k$-ambiguous contract whose principal’s utility is at least $\frac{1}{n-k}$ of the principal’s utility in the unrestricted case.
As a direct corollary, an ambiguous contract of size at most $n-2$ (i.e., just one fewer than the number needed to achieve the optimal utility of the unrestricted case) guarantees at least half of the optimal principal’s utility in the unrestricted case. We show that this is tight.
Theorem (Theorem 5.2). There exists an instance with $n$ actions where no ambiguous contract with strictly less than $n-1$ payment functions yields more than $\frac{1}{2}+O(\frac{1}{n})$ of the optimal principal’s utility in the unrestricted case.
Finally, we provide an additional upper bound on the succinctness gap, which applies to any values of $k$ and $n$. It shows that, in the worst-case, the principal’s utility from a $k$-ambiguous contract can be a factor $\frac{k}{n+k-2}$ smaller than that in the unrestricted case.
Theorem (Theorem 5.3). For every integer $n$ (under some condition) and any $k=1,\ldots,n-2$, there exists an instance with $n$ actions, such that no IC $k$-ambiguous contract yields more than $\frac{k}{n+k-2}$ of the optimal principal’s utility in the unrestricted case.
An interesting implication of this upper bound is that the succinctness gap has to be at least of order $k/n$, showing that a loss proportional to the ratio between the number of contracts in the support of the ambiguous contract and the number of actions is unavoidable.
1.2 Related Work
This paper contributes to the growing literature on algorithmic contract design, and lies at the intersection of several strands of research.
Ambiguous Contracts and Ambiguity-Aversion.
The role of strategic ambiguity as a tool in contracting was first highlighted by Bernheim and Whinston (1998) (albeit in a rather different model and sense than in this work). They consider two-player normal-form games, in which the actions of the two players are partitioned into sets, with the interpretation that a court/some third-party can distinguish between actions from different sets but not between actions that belong to the same set. A contract in their model is a restriction of the allowed actions to a subset of the actions, with a restriction on feasible contracts arising from the partial verifiability of which actions were chosen. A complete contract narrows down the action sets to a single set in each agent’s partition. They find that in static games, under pure Nash equilibria, it is without loss to consider complete contracts. However, they also show that in dynamic settings, where agents choose actions sequentially, optimal contracts may be incomplete.
Our model and notion of ambiguous contracts build on but differ from those in Dütting et al. (2024). Like Dütting et al. (2024), we augment a standard hidden-action principal-agent model with ambiguity and analyze how a principal can optimally introduce ambiguity when interacting with an ambiguity-averse agent with max-min utility. A key innovation in our work is the explicit simplicity constraint: we limit the number of classic contracts that an ambiguous contract may be composed of. Similar to Dütting et al. (2024),we focus on pure actions. Indeed, Dütting et al. (2024) show that, when the agent employs mixed strategies, they can hedge against ambiguity, neutralizing any strategic advantage for the principal.
The notion of ambiguous contracts in Dütting et al. (2024) and this work is inspired by the work of Di Tillio et al. (2017), who introduce an analogous notion in auction design, showing that a seller can exploit ambiguity to achieve higher surplus when selling an object to an ambiguity-averse buyer. They characterize optimal ambiguous mechanisms, but also note their complexity and suggest that this may limit their applicability in practice. To address this, they exhibit a class of simple ambiguous mechanisms (which they call naïve ambiguous mechanisms) and show that these dominate all non-ambiguous mechanisms in terms of the expected revenue they generate. In contrast, we formulate an explicitly (parametric) simplicity constraint, and characterize the optimal ambiguous design for all simplicity levels.
Several other works have considered settings in which the agent holds ambiguous beliefs that the principal can potentially exploit. Bose and Renou (2014) examine mechanism design problems with ambiguous communication devices in which agents have max-min preferences. Beauchêne et al. (2019) and Cheng (2020) consider Bayesian persuasion, in which the sender exploits the ambiguity aversion of the receiver. Bodoh-Creed (2012) examine screening problems with agents who have max-min preferences. Lopomo et al. (2011) explore moral hazard problems with agents who have Bewley preferences. Bose et al. (2006) consider auctions in which the seller and bidders may both be ambiguity-averse.
Simplicity and Robustness in Contract Design.
A recurring theme in contract theory (as well as mechanism design) is the trade-off between simplicity and performance.
Carroll (2015), Carroll and Walton (2022), Kambhampati (2023), and Peng and Tang (2024) examine moral hazard problems in which the principal has ambiguous beliefs about the set of actions the agent can choose from, and establish max-min optimality of linear (a.k.a. commission-based) contracts. Dai and Toikka (2022) examine a principal who writes contracts to shape the actions of a team of agents, with the principal holding ambiguous beliefs about the actions available to the agents, again showing that linear contracts are robustly optimal.
Dütting et al. (2019) examine moral hazard problems in which the principal has ambiguous beliefs about the distribution of outcomes induced by the agent’s actions, once again showing max-min optimality of linear contracts. In addition, Dütting et al. (2019) show (tight) worst-case approximation guarantees of linear contracts, relative to optimal contracts.
Our work adds to this line of work, by formulating an explicit simplicity constraint, and characterizing optimal ambiguous contracts under this constraint.
Algorithmic Approaches to Contracts.
Another adjacent trajectory in the literature takes an algorithmic approach to contracts. An important such direction is the literature on combinatorial contracts. This includes contracting problems with multiple agents (Babaioff et al., 2006, 2012; Dütting et al., 2023; Castiglioni et al., 2023; Dütting et al., 2025), many actions (Dütting et al., 2021a, 2024b; Ezra et al., 2024), or with complex outcome spaces (Dütting et al., 2021b). Another important direction considers typed contract settings, either with multi-dimensional types (Guruganesh et al., 2021; Castiglioni et al., 2021, 2022) or single-dimensional types (Alon et al., 2021, 2023). Castiglioni et al. (2025) have recently established reductions between the two settings.
Our work fits into this literature by taking an algorithmic approach to reveal the structure of optimal succinct ambiguous contracts, while simultaneously shedding light on the computational complexity of finding such contracts.
2 Model and Preliminaries
The basic ingredients of the principal-agent model apply to both classic and ambiguous contracts. We focus here on instances, where all outcomes have non-negative rewards.
Definition 2.1 (Instance).
An instance $\mathcal{I}=(c,r,p)$ of the principal-agent problem with $n$ actions and $m$ outcomes is specified by:
  • A cost vector $c=(c_{1},\ldots,c_{n})$, where for every action $i\in[n]$, $c_{i}\in\mathbb{R}_{+}$ denotes action $i$’s cost. Costs are assumed to be sorted so that $c_{1}\leq c_{2}\leq\ldots\leq c_{n}$.
  • A reward vector $r=(r_{1},\ldots,r_{m})$, where for every outcome $j\in[m]$, $r_{j}\in\mathbb{R}_{+}$ denotes outcome $j$’s reward. Rewards are assumed to be sorted so that $r_{1}\leq r_{2}\leq\ldots\leq r_{m}$.
  • A probability distribution matrix $p=(p_{1},\ldots,p_{n})$, where for every action $i\in[n]$, $p_{i}\in\Delta^{m}$ denotes the probability distribution of action $i$ over outcomes, where $p_{ij}$ denotes the probability of outcome $j$ under action $i$.
The expected reward of action $i$ is given by $R_{i}=\sum_{j=1}^{m}p_{ij}r_{j}$, and the expected welfare of action $i$ is defined as $W_{i}=R_{i}-c_{i}$.
The agent retains the option of not participating. To capture this, we assume throughout that action $1$ is a zero-cost action. We assume that the expected reward of action $1$ is $R_{1}=0$.
It will also be useful to define subinstances. Fix an instance $\mathcal{I}=(c,r,p)$ of the principal-agent problem with $n$ actions and $m$ outcomes. Then for any subset of actions $A\subseteq[n]$ we define subinstance $\mathcal{I}_{A}$ to be the original instance $\mathcal{I}$ restricted to the set of actions $A$. Note that a subsinstance may lose the property of admitting a zero-cost action.
The instance $\mathcal{I}$ is known to both the principal and the agent. The agent’s action is known only by the agent, while realized outcomes are observed by both the principal and the agent.
A payment function $t:[m]\rightarrow\mathbb{R}_{+}$ identifies a payment made by the principal to the agent upon the realization of an outcome, with the payment $t(j)$ made in response to outcome $j$ typically denoted by $t_{j}$. Note that the payment is a function of the outcome rather than the actions, since the principal cannot observe the agent’s action, only the outcome. As standard, we impose the limited liability assumption, requiring that the payments are non-negative.
2.1 Classic Contracts
We first describe the classic setting, in which a contract, denoted by $\langle t,i\rangle$, is a payment function $t$ and a recommended action $i\in[n]$.
The interpretation is that the principal posts a contract, the agent observes the contract and chooses an action and bears the attendant cost, an outcome is drawn from the distribution over outcomes induced by that action, and the principal makes the payment to the agent specified by the contract. The inclusion of a recommended action in the contract allows us to capture the common presumption that the agent “breaks ties in favor of the principal.”
More precisely, an agent who chooses action $i^{\prime}$ when facing a contract $\langle t,i\rangle$ garners expected utility
$U_{A}({i^{\prime}}\mid{t})~{}~{}=~{}~{}\sum_{j=1}^{m}p_{i^{\prime}j}t_{j}-c_{i^{\prime}}~{}~{}=~{}~{}T_{i^{\prime}}(t)-c_{i^{\prime}},$
given by the difference between the expected payment $T_{i^{\prime}}(t)$ and the cost $c_{i^{\prime}}$. The resulting principal’s utility is $U_{P}({i^{\prime}}\mid{t})=R_{i^{\prime}}-T_{i^{\prime}}(t)$, defined as the difference between expected reward $R_{i^{\prime}}$ and payment $T_{i^{\prime}}(t)$.
A central definition is the following definition of an incentive compatible (or IC) contract, and the associated notion of “implementing an action.”
Definition 2.2 (IC contract).
A contract $\langle t,i\rangle$ is incentive compatible (IC) if $i\in\operatorname*{arg\,max}_{i^{\prime}\in[n]}U_{A}({i^{\prime}}\mid{t}),$ in which case we say that contract $\langle t,i\rangle$ implements action $i$.
Because payments are non-negative and action $1$ has zero cost, incentive compatibility ensures that the agent secures an expected utility of at least zero, and hence implies individual rationality.
We assume the agent follows the recommendation of an incentive-compatible contract. If the principal posts the IC contract $\langle t,i\rangle$, the payoffs to the principal and agent are then
$U_{P}(\langle t,i\rangle)~{}=~{}U_{P}({i}\mid{t})=R_{i}-T_{i}(t)\quad\text{and}\quad U_{A}(\langle t,i\rangle)~{}=~{}U_{A}({i}\mid{t})=T_{i}(t)-c_{i}.$
It is without loss of generality to restrict the principal to incentive compatible contracts. The idea is that an agent facing contract $\langle t,i\rangle$ will choose an action that maximizes her expected utility given $t$, and hence the principal might as well name such an action in the contract.
Optimal Classic Contracts.
An optimal classic contract maximizes the principal’s expected utility. It’s well known that optimal contracts can be computed in polynomial time by solving one linear program (LP) for each action (e.g., Grossman and Hart, 1983; Dütting et al., 2019). The LP for action $i$ minimizes the expected payment that is required to incentivize the agent to take action $i$, and thus identifies the contract that maximizes the principal’s utility from action $i$.
The LP and its dual are given in Figure 1. We refer to these as MIN-PAY-LP$(i)$ and DUAL-MIN-PAY-LP$(i)$. The variables in the primal LP are the payments $\{t_{j}\}_{j\in[m]}$. The variables in the dual are $\{\lambda_{i^{\prime}}\}_{i^{\prime}\in[n]\setminus\{i\}}.$
$\min\quad$ $\sum_{j}p_{ij}t_{j}$
s.t. $\sum_{j}p_{ij}t_{j}-c_{i}\geq\sum_{j}p_{i^{\prime}j}t_{j}-c_{i^{\prime}}$ $\forall i^{\prime}\neq i$
$t_{j}\geq 0$ $\forall j$
(a) MIN-PAY-LP$(i)$
$\max\quad$ $\sum_{i^{\prime}\neq i}\lambda_{i^{\prime}}(c_{i}-c_{i^{\prime}})$
s.t. $\sum_{i^{\prime}\neq i}\lambda_{i^{\prime}}(p_{ij}-p_{i^{\prime}j})\leq p_{ij}$ $\forall j$
$\lambda_{i^{\prime}}\geq 0$ $\forall i^{\prime}\neq i$
(b) DUAL-MIN-PAY-LP$(i)$
Figure 1: The minimum payment LP for action $i$ and its dual.
We note that MIN-PAY-LP$(i)$ may be feasible or infeasible. It is feasible precisely when action $i$ can be implemented by some classic contract. Whenever it is feasible, we refer to any contract $\langle t,i\rangle$ that optimally solves the LP, as a min-pay contract for action $i$.
Remark 2.1 (Non-zero payments).
Standard LP considerations (e.g. Matous̆ek and Gärtner, 2006) imply that the min-pay contract has at most $n-1$ non-zero payments.
Implementability by Classic Contracts.
The LP formulation can also be leveraged to yield a characterization of actions that are implementable with classic contracts. This is cast in the following proposition.
Proposition 2.1 (Hermalin and Katz (1991)).
Action $i\in[n]$ is implementable with a classic contract if and only if there does not exist a convex combination $\lambda_{i^{\prime}}\in[0,1]$ of the actions $i^{\prime}\neq i$ that yields the same distribution over rewards $\sum_{i^{\prime}\neq i}\lambda_{i^{\prime}}p_{i^{\prime}j}=p_{ij}$ for all $j$ but at a strictly lower cost $\sum_{i^{\prime}}\lambda_{i^{\prime}}c_{i^{\prime}}<c_{i}$.
2.2 Ambiguous Contracts
We next define the notion of ambiguous contracts, as introduced in Dütting et al. (2024). The basic situation is similar to that in the classic setting, except that now the principal posts a collection of payment functions.
Definition 2.3 (Ambiguous contract).
An ambiguous contract $\langle\tau,i\rangle$ is a collection of payment functions $\tau=\{t^{1},t^{2},\ldots\}$ and a recommended action $i$. If $t\in\tau$, we say $t$ is in the support of $\tau$.
It will be convenient to allow a collection of payment functions $\tau$ to be a multiset. This is generally without loss, and avoids having to explicitly deal with duplicates.
The agent is assumed to be ambiguity-averse, employing a max-min utility (Schmeidler, 1989; Gilboa and Schmeidler, 1993). Formally, we define the agent’s utility for action $i$ given a collection of payment functions $\tau$ as
$U_{A}(i\mid\tau)=\min_{t\in\tau}U_{A}(i\mid t).$
The agent’s best response to a collection of payment functions is then the action $i$ that maximizes the agent’s (minimum) utility.
Definition 2.4 (Best response).
Action $i$ is a best response to a collection of payment functions $\tau$ if it holds that
$U_{A}(i\mid\tau)\geq U_{A}(i^{\prime}\mid\tau)\quad\text{for all$i^{\prime}\neq i$.}$
Following Dütting et al. (2024) we also insistent on “consistency,” defined as follows:
Definition 2.5 (Consistency).
A collection of payment functions $\tau$ is consistent w.r.t. action $i$ if
$U_{P}({i}\mid{t^{\ell}})=U_{P}({i}\mid{t^{\ell^{\prime}}})\quad\text{for all $t^{\ell},t^{\ell^{\prime}}\in\tau$.}$ (1)
In other words, consistency requires that, under the given target action $i$, the principal’s utility is the same for all payment functions in the collection of payment functions $\tau$. Without this assumption, the principal’s “threat” to use any contract in the support of the ambiguous contract is implausible.
Definition 2.6 (IC ambiguous contract).
An ambiguous contract $\langle\tau,i\rangle$ is incentive compatible (IC) if (i) action $i$ is a best response to $\tau$, and (ii) $\tau$ is consistent with respect to action $i$. In this case, we say that $\langle\tau,i\rangle$ implements action $i$.
If the principal posts the IC ambiguous contract $\langle\tau,i\rangle=\langle\{t^{1},t^{2},\ldots\},i\rangle$, then, by consistency, the induced payment $T_{i}(\tau)$ can be defined as
$T_{i}(\tau)=T_{i}(t^{\ell})=T_{i}(t^{\ell^{\prime}})\quad\mbox{for every}\quad\ell,\ell^{\prime},$
and since the agent’s utility is $U_{A}({i}\mid{t})=T_{i}(t)-c_{i}$, it holds that
$U_{A}({i}\mid{t^{\ell}})=U_{A}({i}\mid{t^{\ell^{\prime}}})\quad\mbox{for every}\quad\ell,\ell^{\prime}.$
Consequently, the principal’s and agent’s expected utilities satisfy the following for any $t\in\tau$:
$U_{P}(\langle\tau,i\rangle)=U_{P}({i}\mid{t})~{}=~{}R_{i}-T_{i}(\tau)\quad\text{and}\quad U_{A}(\langle\tau,i\rangle)=U_{A}({i}\mid{t})~{}=~{}T_{i}(\tau)-c_{i}.$
While the choice to insist on consistency may seem restrictive at first glance, it turns out to be without loss of generality:
Lemma 2.1 (Dütting et al. (2024)).
Suppose action $i$ is a best response to a collection of payment functions $\tau$. Then there exists a (consistent) incentive compatible ambiguous contract $\langle\tau^{\prime},i\rangle$ from which the principal obtains expected payoff at least $\max_{t\in\tau}U_{P}({i}\mid{t})$.
As an implication, we can equivalently view the principal as selecting the contract to be implemented, or as max-min preferences over the outcome of a selection by a third party.
Optimal Ambiguous Contracts.
An optimal ambiguous contract maximizes the principal’s expected utility. In Dütting et al. (2024), it is shown that optimal ambiguous contract take a certain form. We first introduce the simplicity notion of a single-outcome payment contract.
Definition 2.7 (SOP payment function).
A payment function $t=(t_{1},\ldots,t_{m})$ is a single-outcome payment function if there exists $j\in[m]$ such that $t_{j}>0$, and for any outcome $j^{\prime}\neq j$, $t_{j^{\prime}}=0$.
An ambiguous contract is a single-outcome payment contract if all of its payment functions have this property.
The characterization result now states that it is without loss of generality to consider ambiguous contracts that are composed of SOP payment functions, and that at most $\min\{m,n-1\}$ SOP payment functions suffice.
Theorem 2.1 (Optimal ambiguous contracts, Dütting et al. (2024)).
For every IC ambiguous contract $\langle\tau,i\rangle$, there exists an IC ambiguous contract $\langle\tau^{\prime},i\rangle$, containing at most $\min\{m,n-1\}$ payment functions, such that:
  1. 1.
    $T_{i}(\tau^{\prime})=T_{i}(\tau)$ (i.e., both contracts have the same expected payment and hence same expected payoff to the principal).
  2. 2.
    For every $t^{\prime}\in\tau^{\prime}$, $t^{\prime}$ is an SOP payment function.
Dütting et al. (2024) also give the following “closed form” characterization of the optimal ambiguous SOP contract.
Proposition 2.2 (Closed form optimal ambiguous contract, Dütting et al. (2024)).
Suppose action $i\in[n]$ is implementable by an ambiguous contract. Let $A=\{i^{\prime}\neq i\mid p_{i^{\prime}}\neq p_{i}\}$. If $A=\emptyset$, then $c_{i}=0$, and the IC contract $\langle\{(0,\ldots,0)\},i\rangle$ is optimal for action $i$. Otherwise, for each $i^{\prime}\in A$ let $j(i^{\prime})$ be an outcome such that $p_{ij(i^{\prime})}/p_{i^{\prime}j(i^{\prime})}$ is maximal. Let
$T=\max_{i^{\prime}\in A}\left\{\min\left\{x\geq 0\;\middle|\;p_{ij(i^{\prime})}\cdot\frac{x}{p_{ij(i^{\prime})}}-c_{i}\geq p_{i^{\prime}j(i^{\prime})}\cdot\frac{x}{p_{ij(i^{\prime})}}-c_{i^{\prime}}\right\}\right\}.$
For each $i^{\prime}\in A$, let $t^{i^{\prime}}_{j(i^{\prime})}=T/p_{ij(i^{\prime})}$ and $t^{i^{\prime}}_{j^{\prime}}=0$ for $j^{\prime}\neq j(i^{\prime})$. Then the IC contract $\langle\tau,i\rangle=\langle\{t^{i^{\prime}}\mid i^{\prime}\in A\},i\rangle$ is optimal for action $i$.
Implementability by Ambiguous Contracts.
Dütting et al. (2024) also obtain the following characterization of actions that can be implemented with an ambiguous contract.
Proposition 2.3 (Dütting et al. (2024)).
Action $i\in[n]$ is implementable with an ambiguous contract if and only if there is no other action $i^{\prime}\neq i$ such that $p_{i^{\prime}}=p_{i}$ and $c_{i^{\prime}}<c_{i}$.
Compared to the conditions for implementability by a classic contract (Proposition 2.1), it is clear that the conditions for implementing an action with an ambiguous contracts are more permissive.
3 Succinct Ambiguous Contracts: Structure and Algorithms
The focus of this work is on what we call succinct ambiguous contracts. Informally, these are contracts that are composed of at most $k$ payment functions. For technical reasons, it will be convenient to define them as ambiguous contracts that contain precisely $k$ payment functions.
This is without loss, because we treat collections of payment functions as multisets. So we can always “fill up” a collection of payment functions that contains fewer than $k$ payment functions. [^1] We note that, one may wish to identify the smallest $k^{\prime}\leq k$ that achieves a given utility. In this case, one can run our algorithms for each $k^{\prime}\leq k$ and determine the smallest $k^{\prime}$ satisfying the utility requirement.
Definition 3.1 ($k$-ambiguous contract).
For $k\in\mathbb{N}$, a $k$-ambiguous contract $\langle\tau,i\rangle=\langle\{t^{1},\ldots,t^{k}\},i\rangle$ is a collection of $k$ payment functions and a recommended action.
We know from Theorem 2.1 that for $k\geq\min\{n-1,m\}$ there is no difference between $k$-ambiguous contracts and general ambiguous contracts. Moreover, in this unrestricted case, Theorem 2.1 shows that the optimal IC ambiguous contract is composed of SOP payment functions. Proposition 2.2 further shows that these SOP contracts concentrate payments on the maximum likelihood-ratio outcomes. Such contracts are known to be optimal for settings with two actions (e.g., Laffont and Martimort, 2009, Chapter 4.5.1.) (also see (Dütting et al., 2019, Proposition 5)).
Our main result in this section (Theorem 3.1) is a generalization of these results, providing a structural characterization of optimal $k$-ambiguous contracts for any $k\geq 1$. We also deduce from it a characterization of actions that can be implemented with $k$-ambiguous contracts (Proposition 3.1), again generalizing the characterization that applies in the unrestricted case.
Our proofs in this section provide alternative proofs of the key results of Dütting et al. (2024) for the unrestricted case; and show that optimal $k$-ambiguous contracts are governed by a “divide and conquer” principle (more on this below).
3.1 Optimal $k$-Ambiguous Contracts
We first state our main characterization result for $k$-ambiguous contracts. To formally state the result it will be useful to recall a few facts about optimal classic contracts (see Section 2.1). Most notably, that for any action $i$ that is implementable by a classic contract, we can find an optimal (classic) contract $\langle t,i\rangle$ for implementing it by solving an LP (namely, MIN-PAY-LP$(i)$ in Figure 1). Recall that we refer to such contracts as min-pay contracts.
Our characterization says that an optimal $k$-ambiguous contract $\langle\tau,i^{\star}\rangle$ with $\tau=\{t^{1},\ldots,t^{k}\}$ is essentially a collection of $k$ min-pay contracts $\langle t^{\ell},i^{\star}\rangle$ obtained by solving MIN-PAY-LP$(i^{\star})$ for $k$ subinstances, each obtained by restricting to the set of actions to $\{i^{\star}\}\cup S^{\ell}$ such that the sets $\{S^{1},\ldots,S^{k}\}$ form a partition of the actions other than $i^{\star}$. It is only “essentially” because the contracts $t^{1},\ldots,t^{k}$ are not the resulting min-pay contracts, but rather the resulting min-pay contracts shifted by adding some amount $c^{\ell}\geq 0$ to all payments, so as to equalize the expected payment from each of these contracts for action $i^{\star}$. We refer to such contracts as shifted min-pay contracts.
One interpretation of this is that optimal $k$-ambiguous contracts can be found by a “divide and conquer” approach: First break the problem into smaller problems that we know how to solve. Then combine the solutions for the smaller problems into a solution to the original problem.
Theorem 3.1 (Optimal $k$-ambiguous contracts).
For any $k$, there is an optimal IC $k$-ambiguous contract $\langle\tau,i^{\star}\rangle$ with $\tau=\{t^{1},\ldots,t^{k}\}$ that takes the following form:
  • There is a partition of the actions $[n]\setminus\{i^{\star}\}$ into $k$ sets $\{S^{1},\ldots,S^{k}\}$.
  • For each $\ell\in[k]$, contract $t^{\ell}$ is a shifted min-pay contract for action $i^{\star}$ for the subinstance $\mathcal{I}_{\{i^{\star}\}\cup S^{\ell}}$, obtained by restricting the original instance to actions $\{i^{\star}\}\cup S^{\ell}$.
Theorem 3.1 is a generalization of Theorem 2.1 since if the ambiguous contract has no size limit, the partition of $[n]\setminus\{i^{\star}\}$ has at most $n-1$ parts, each of size $1$. Moreover, to “protect” action $i^{\star}$ against a set of size $1$, it is without loss of generality to only consider SOP contracts (e.g. Laffont and Martimort, 2009, Chapter 4.5.1) (also see (Dütting et al., 2019, Proposition 5)). Moreover, in that case, the upper bound on $|\tau|$ of $n-1$ can be strengthened to $\max\{n-1,m\}$ because if two SOP contracts in the support of $\tau$ pay for the same outcome, one of them is redundant.
Proof Outline.
To prove Theorem 3.1 we proceed as follows. First, in Section 3.2 we define the concept of protecting a target action against a partition of the actions different from the target action into $k$ sets. Then, in Section 3.3, we describe a balancing routine that combines a collection of payment functions that protect a target action against a partition of the other actions into $k$ sets into an IC $k$-ambiguous contract. Afterwards, in Section 3.4, we show how to find an optimal IC $k$-ambiguous contract for the target action and a fixed partition of the other actions into $k$ sets. Finally, in Section 3.5, we show that by iterating over all actions and partitions, we obtain an optimal ambiguous contract.
3.2 Protecting Actions against Sets of Actions and Partitions of the Actions
Our proof of Theorem 3.1 revolves around the idea of protecting a target action $i$ against a subset of the actions other than $i$, and more generally a partition of the actions other than $i$ into $k$ sets. We first define the notion of protecting an action against a set of actions.
Definition 3.2.
A payment function $t$ is said to protect action $i\in[n]$ against a set of actions $S\subseteq[n]\setminus\{i\}$ if $U_{A}(i^{\prime}\mid t)\leq U_{A}(i\mid t)$ for all $i^{\prime}\in S$.
We next define what it means for a set of $k$ payment functions to protect an action against a partition of size $k$. Let $\mathcal{T}_{k}$ be the set of all collections of $k$ payment functions. Let $\mathcal{S}_{k,i}$ denote all partitions of $[n]\setminus\{i\}$ into $k$ sets $\{S^{1},\ldots,S^{k}\}$. That is, for each $\ell\in[k]$, $S^{\ell}\subseteq[n]\setminus\{i\}.$ Moreover, $S^{\ell}\cap S^{\ell^{\prime}}=\emptyset$ for all $\ell,\ell^{\prime}\in[k]$ with $\ell\neq\ell^{\prime}$ and $\bigcup_{\ell=1}^{k}S^{\ell}=[n]\setminus\{i\}.$
Definition 3.3.
A collection of payment functions $\{t^{1},\ldots,t^{k}\}\in\mathcal{T}_{k}$ is said to protect action $i\in[n]$ against a partition $\{S^{1},\ldots,S^{k}\}\in\mathcal{S}_{k,i}$ if, for each $j\in[k]$, $t^{j}$ protects action $i$ against $S^{j}$
Note that Definition 3.3 does not yet imply that the ambiguous contract $\langle\{t^{1},\ldots,t^{k}\},i\rangle$ is IC. This is because different payment functions in the support of $\tau$ may entail different expected payments (and hence principal and agent utilities) for action $i$.
3.3 A Balancing Routine that Ensures Consistency
As a first step, we show how to turn a collection of payment functions $\{t^{1},\ldots,t^{k}\}$ that protects action $i$ against a partition $\{S^{1},\ldots,S^{k}\}\in\mathcal{S}_{k,i}$ into an IC ambiguous contract $\langle\tau,i\rangle$ of support size $k$. Namely, we show how to turn it into a consistent collection of payment functions for which action $i$ is a best response.
To this end, consider Algorithm 1. The algorithm is given a target action $i$ and a collection of payment functions $\{t^{1},\ldots,t^{k}\}$. The algorithm first computes the expected payment $T^{\ell}_{i}$ for action $i$ of each payment function $t^{\ell}\in\{t^{1},\ldots,t^{k}\}$, it then sets $\theta:=\max\{T_{i}^{1},\ldots,T_{i}^{k}\}$, and raises each component of the payment function $t^{\ell}$ by the same amount $c^{\ell}=\theta-T_{i}^{\ell}$.
Input: An action $i$, and a collection of payment functions $\{t^{1},t^{2},\ldots t^{k}\}$
Output: An ambiguous contract $\langle\tau,i\rangle$
$\theta$leftarrow $\max\{T_{i}^{1},...T_{i}^{k}\}$ where, for each $j\in[k]$, $T_{i}^{j}=\sum_{\ell\in[m]}p_{i\ell}t^{j}_{\ell}$;
$\tau=\emptyset$;
for $j\in[k]$  do
      $\bar{t}^{j}$leftarrow $t^{j}+(\theta-T^{j}_{i})\cdot\vec{1}$ (for consistency);
      $\tau$leftarrow $\tau\cup\{\bar{t}^{j}\}$;
      end for
return $\langle\tau,i\rangle$
Algorithm 1 A Balancing Routine that Ensures Consistency
The following lemma shows that, when Algorithm 1 is fed with a collection of payment functions $\{t^{1},\ldots,t^{k}\}$ that protect action $i$ against a partition $\mathcal{S}=(S^{1},\ldots,S^{k})$ of the actions other than $i$, then it returns an (ambiguous) IC contract $\langle\tau,i\rangle.$ The proof appears in Appendix A.
Lemma 3.1.
Fix an action $i\in[n]$ and a collection of payment functions $\{t^{1},\ldots,t^{k}\}$ that protects action $i$ against a partition $\{S^{1},\ldots,S^{k}\}\in\mathcal{S}_{k,i}$. Then Algorithm 1 returns an IC ambiguous contract $\langle\tau,i\rangle$ with $\tau=\{\bar{t}^{1},\ldots,\bar{t}^{k}\}$ satisfying $U_{P}({i}\mid{\bar{t}^{1}})=\ldots=U_{P}({i}\mid{\bar{t}^{k}})=\min_{j\in[k]}\{U_{P}({i}\mid{t^{j}})\}$.
3.4 Finding IC Ambiguous Contracts that are Optimal for a Given Partition
As a next step, we show how to find an IC $k$-ambiguous contracts that is optimal for a given target action and fixed partition of the other actions, in the following sense.
Definition 3.4.
Given an action $i\in[n]$, an IC $k$-ambiguous contract $\langle\tau,i\rangle$ is said to be optimal for action $i$ with respect to a partition $\mathcal{S}\in\mathcal{S}_{k,i}$ if (i) $\tau$ protects action $i$ against $\mathcal{S}$, and (ii) for any other IC $k$-ambiguous contract $\langle\tau,i\rangle$ such that $\hat{\tau}$ protects action $i$ against $\mathcal{S}$, it holds that $U_{P}({i}\mid{\hat{\tau}})\leq U_{P}({i}\mid{\tau})$.
Note that Definition 3.4 only constitutes a necessary condition for $\langle\tau,i\rangle$ to be an optimal IC $k$-ambiguous contract. This is because, if $\langle\tau,i\rangle$ is an optimal IC $k$-ambiguous contract, then it is optimal for action $i$ with respect to some partition $\mathcal{S}\in\mathcal{S}_{k,i}$.
[…]
Input: An action $i\in[n]$, a partition $\mathcal{S}=\{S^{1},\ldots,S^{k}\}$ of $[n]\setminus\{i\}$
Output: An IC $k$-ambiguous contract $\langle\tau,i\rangle$ that is optimal for action $i$ with respect to $\mathcal{S}$, or Null if there is no $\tau\in\mathcal{T}_{k}$ that protects action $i$ against $\mathcal{S}$
if $i$ is not implementable with respect to $S^{j}\cup\{i\}$ for some $j\in[k]$  then
      return Null
      else
           For all $j\in[k]$:  $t^{j}$leftarrow the solution to MIN-PAY-LP$(i)$ on subinstance $\mathcal{I}_{\{i\}\cup S^{j}}$ ;
           $\langle\tau,i\rangle$leftarrow Algorithm 1 ($i$, $\{t^{1},\ldots,t^{k}\}$);
           end if
return $\langle\tau,i\rangle$
Algorithm 2 Optimal IC $k$-Ambiguous Contract for a Given Partition
The following lemma shows that Algorithm 2 indeed finds an optimal IC $k$-ambiguous contract for action $i$ and a fixed partition $\mathcal{S}=\{S^{1},\ldots,S^{k}\}$ of the actions other than $i$; and shows that the resulting ambiguous contract consists of shifted min-pay contracts. The proof is in Appendix A.
Lemma 3.2.
Given an action $i$ and a partition $\mathcal{S}=\{S^{1},\ldots,S^{k}\}\in\mathcal{S}_{k,i}$, Algorithm 2 returns an optimal IC $k$-ambiguous contract for action $i$ with respect to $\mathcal{S}$, or Null if there is no $\tau\in\mathcal{T}_{k}$ that protects action $i$ against $\mathcal{S}$. Moreover, in the case where the algorithm returns an ambiguous contract $\langle\tau,i\rangle$, each payment function in the support of $\tau$ corresponds to a shifted min-pay contract.
3.5 From Optimality for a Given Partition to Optimal $k$-Ambiguous Contracts
As a final step, we now show that, for any fixed target action $i$, we can find an optimal IC $k$-ambiguous contract for that action, or decide that none exists. This is achieved by Algorithm 3, which applies Algorithm 2 to all possible partitions $\mathcal{S}\in\mathcal{S}_{k,i}$ of the actions other than $i$. For each partition $\mathcal{S}$, the algorithm thus finds an IC $k$-ambiguous contract $\langle\tau,i\rangle$ that is optimal with respect to $\mathcal{S}$ (or that it is impossible to protect action $i$ against this partition). It then return the IC $k$-ambiguous contract $\langle\tau^{\mathcal{S}},i\rangle$ that minimizes the expected payment required to implement action $i$ (if one exists). In general, this algorithm is an exponential-time algorithm, as it iterates over all partitions in $\mathcal{S}_{k,i}$. However, with small adjustments, there are regimes of $k$ where its running time is poly$(n,m)$. See Remark 3.1.
The following lemma shows the correctness of Algorithm 3. It shows that the optimal $k$-ambiguous contract that implements a given target action $i$ (if one exists) is of the form that is claimed to be optimal overall in Theorem 3.1.
Lemma 3.3.
Given an action $i$, Algorithm 3 finds an optimal IC $k$-ambiguous contract $\langle\tau,i\rangle$, or decides that no such contract exists. In the case, where action $i$ is implementable, the resulting $k$-ambiguous contract consists of shifted min-pay contracts.
Proof.
First, consider the case where action $i$ is implementable by a $k$-ambiguous contract. That is, there exists an IC ambiguous contract of size $k$ that implements action $i$. Let $\langle\hat{\tau}=\{\hat{t}^{1},\ldots,\hat{t}^{k}\},i\rangle$ be an arbitrary such contract. It suffices to show that Algorithm 3 returns an IC $k$-ambiguous contract $\langle\tau^{\mathcal{S}},i\rangle$ such that $U_{P}(\langle\tau^{\mathcal{S}},i\rangle)\geq U_{P}(\langle\hat{\tau},i\rangle)$. Let $\hat{\mathcal{S}}=\{\hat{S}^{1},\ldots,\hat{S}^{k}\}$ be a partition of $[n]\setminus\{i\}$ defined as follows:
$\hat{S}^{1}=\{i^{\prime}\in[n]\setminus\{i\}\mid U_{A}({i^{\prime}}\mid{\hat{t^{1}}})\leq U_{A}({i}\mid{\hat{t}^{1}})\}$
$\hat{S}^{2}=\{i^{\prime}\in\left([n]\setminus\{i\}\right)\setminus\hat{S}^{1}\mid U_{A}({i^{\prime}}\mid{\hat{t}^{2}})\leq U_{A}({i}\mid{\hat{t}^{2}})\}$
$\vdots$
$\hat{S}^{k}=\left\{i^{\prime}\in\left([n]\setminus\{i\}\right)\setminus\bigcup_{\ell\in[k-1]}\hat{S}^{\ell}\;\bigg{|}\;U_{A}({i^{\prime}}\mid{\hat{t}^{k}})\leq U_{A}({i}\mid{\hat{t}^{k}})\right\}$
By design, the collection of payment functions $\hat{\tau}$ protects action $i$ against partition $\hat{\mathcal{S}}$. Since $\hat{\mathcal{S}}$ is in $\mathcal{S}_{k,i}$, Algorithm 3 iterates over it. It follows that $\mathcal{C}\neq\emptyset$ (since it contains at least $\langle\tau^{\hat{\mathcal{S}}},i\rangle$). Moreover, the contract $\langle\tau^{\mathcal{S}},i\rangle$ returned by Algorithm 3 has principal utility at least as high as $\langle\tau^{\hat{\mathcal{S}}},i\rangle$, which, in turn, has principal utility at least as high as $\langle\hat{\tau},i\rangle$ by Lemma 3.2.
Next, consider the case where action $i$ is not implementable by an ambiguous contract of size $k$. In this case, no partition of $[n]\setminus\{i\}$ to $k$ sets would result in an output that is different from Null (when applying Algorithm 2 to action $i$ and the given partition). Thus $\mathcal{C}=\emptyset$ and Algorithm 3 returns Null. ∎
Input: An action $i$
Output: An optimal IC ambiguous contract $\langle\tau,i\rangle$ with $\tau=\{t^{1},\ldots,t^{k}\}$ for implementing action $i$, or Null if the action is not implementable
$\mathcal{C}\leftarrow$emptyset;
for $\mathcal{S}\in\mathcal{S}_{k,i}$ do
     output $\leftarrow$ Algorithm 2 ($i$, $\mathcal{S}$);
      if output $\neq$ Null  then
           $\langle\tau^{\mathcal{S}},i\rangle$leftarrow output;
           $\mathcal{C}\leftarrow\mathcal{C}\cup\{\langle\tau^{\mathcal{S}},i\rangle\}$;
           end if
           end for
          return Null if $C=\emptyset$, or $\langle\tau^{\mathcal{S}},i\rangle$ where $\mathcal{S}\in\operatorname*{arg\,min}_{\mathcal{S}:\;\langle\tau^{\mathcal{S}},i\rangle\in\mathcal{C}}T_{i}(\tau^{\mathcal{S}})$ otherwise;
Algorithm 3 Optimal $k$-Ambiguous Contract for a Given Action
In order to find an optimal ambiguous contract of size at most $k$, we can apply Algorithm 3 to all actions and choose the action that leads to the highest principal utility.
Proof of Theorem 3.1.
The best $k$-ambiguous contract overall implements some action $i^{\star}$. Since we iterate over all actions, the statement of the theorem follows from Lemma 3.3. ∎
Remark 3.1 (Running time).
[…]
3.6 Implementability under $k$-Ambiguous Contracts
The structural properties revealed above lend themselves to a characterization of implementability by a $k$-ambiguous contract. This is cast in the following proposition, which we prove in Appendix A.
Proposition 3.1.
An action $i$ in an instance $\mathcal{I}$ is implementable by a $k$-ambiguous contract if and only if there exists a partition $\mathcal{S}$ of $[n]\setminus\{i\}$ to $k$ sets, such that for every set $S\in\mathcal{S}$, action $i$ is implementable by a classic contract in subinstance $\mathcal{I}_{\{i\}\cup S}$. That is, for each $S\in\mathcal{S}$ there does not exist a convex combination $\lambda_{i^{\prime}}\in[0,1]$ of the actions $i^{\prime}\in S$ that yields the same distribution over rewards $\sum_{i^{\prime}\in S}\lambda_{i^{\prime}}p_{i^{\prime}j}=p_{ij}$ for all $j$ but at a strictly lower cost $\sum_{i^{\prime}\in S}\lambda_{i^{\prime}}c_{i^{\prime}}<c_{i}$.
Proposition 3.1 is trivially a generalization of Proposition 2.1, simply because Proposition 3.1 instantiated with $k=1$ coincides with Proposition 2.1. It is also a generalization of Proposition 2.3 since if the ambiguous contract has no size limit, the partition of $[n]\setminus\{i\}$ has at most $n-1$ parts, of size $1$, and a convex combination of a set of one action $\{i^{\prime}\}$ is always equal to the action itself.
As a corollary of Proposition 3.1 we obtain that, if an action is implementable by a $k$-ambiguous contract, then it is also implementable by any $k^{\prime}$-ambiguous contract, for all $k^{\prime}\geq k$. Hence, growing $k$ from $1$ to $n-1$ amounts to a gradual relaxation of the implementability constraints, with classic contracts and unrestricted ambiguous contracts representing the two extremes of this hierarchy.
4 Optimal Succinct Ambiguous Contract Problem: Hardness
Section 3 presents an algorithm for finding an optimal $k$-ambiguous contract. For each action $i\in[n]$, the algorithm iterates over all (potentially exponentially many) partitions of $[n]\setminus\{i\}$ into $k$ sets, reducing the problem to finding the optimal partition among them. This process is inherently complex computationally, and as discussed in Remark 3.1, leads to a super-polynomial-in-$n$ running time unless $k=n-O(1)$.
This raises the natural question of whether a more efficient algorithm could bypass this complexity. In this section, we address this question by establishing the intrinsic hardness of the problem. We do so by reducing the well-known Makespan Minimization problem (e.g., Garey and Johnson, 1979, Section A5.2) into the optimal succinct ambiguous contract problem. Thus, the optimal $k$-ambiguous contract problem is NP-hard for any $k$ for which finding a partition of $n$ numbers into $k$ sets that minimizes the makespan is NP-hard. This includes any constant $k$ or any $k=\beta\cdot n$ for $\beta\in(0,1)$ (see Appendix B for details).
The following defines the well-known Makespan Minimization problem.
Definition 4.1 (Makespan Minimization).
[…]
The following theorem, our main result in this section, establishes the hardness of the optimal $k$-ambiguous contract problem.
Theorem 4.1.
The $k(n)$-ambiguous contract problem is NP-hard for all functions $k:[n]\rightarrow[n]$ such that $(n,k(n))$-Makespan Minimization is NP-hard. In particular, this is true for any constant $k$ and any $k(n)=\beta\cdot n$ for $\beta\in(0,1)$.
To prove Theorem 4.1, we establish the following reduction:
Proposition 4.1.
There is a poly-time algorithm for mapping instances $\mathcal{J}_{n,k}$ of $(n,k)$-Makespan Minimization to instances $\mathcal{I}_{n+2}$ of the $k$-ambiguous contract problem, and a poly-time algorithm for turning an optimal solution to the $k$-ambiguous contract problem in $\mathcal{I}_{n+2}$ into an optimal solution to the $(n,k)$-Makespan Minimization in $\mathcal{J}_{n,k}$.
Given an instance $\mathcal{J}_{n,k}$ of $(n,k)$-Makespan Minimization with values $a_{1},\ldots a_{n}$, we construct an instance $\mathcal{I}_{n+2}=\mathcal{I}(a_{1},\ldots,a_{n})$ of the $k$-ambiguous contract problem, with $n+2$ actions and $n+1$ outcomes, as depicted in Figure 2.
Next we show that, given an optimal solution to the $k$-ambiguous contract problem in $\mathcal{I}(a_{1},\ldots,a_{n})$, we can construct an optimal solution to $(n,k)$-Makespan Minimization in $\mathcal{J}_{n,k}$.
   rewards:    $r_{0}=0$    $r_{1}=A$    $r_{2}=A$    $\dots$    $r_{n}=A$    costs
action 00: $1$ 00 00 $\dots$ 00 $c_{0}=0$
action $1$: $\frac{1}{n}$ 00 $\frac{1}{n}$dots $\frac{1}{n}$ $c_{1}=c_{n+1}-\frac{1}{n}\cdot a_{1}$
action $2$: $\frac{1}{n}$ $\frac{1}{n}$ 00 $\dots$ $\frac{1}{n}$ $c_{2}=c_{n+1}-\frac{1}{n}\cdot a_{2}$
$\vdots$ $\vdots$ $\vdots$ $\vdots$ $\ddots$ $\vdots$ $\vdots$
action $n$: $\frac{1}{n}$ $\frac{1}{n}$ $\frac{1}{n}$dots 00 $c_{n}=c_{n+1}-\frac{1}{n}\cdot a_{n}$
action $n+1$: 00 $\frac{1}{n}$ $\frac{1}{n}$dots $\frac{1}{n}$ $c_{n+1}=\frac{1}{n}\max_{i\in[n]}a_{i}$
Figure 2: Instance $\mathcal{I}(a_{1},\ldots,a_{n})$ used in the proof of Theorem 4.1.
We first show that the optimal $k$-ambiguous contract implements action $n+1$.
Proposition 4.2.
In the instance $\mathcal{I}(a_{1},\ldots,a_{n})$, the optimal $k$-ambiguous contract implements action $n+1$.
Proof.
To prove the proposition, we construct a classic contract that implements action $n+1$ with a principal’s utility of $\frac{n-1}{n}A$. This shows that the principal’s utility in the optimal k-ambiguous contract is at least $\frac{n-1}{n}A$. On the other hand, it is easy to verify that for any action $i\in[n]$, $R_{i}=\frac{n-1}{n}A$, while for action 00, $R_{0}=0$. The proof then follows since the principal’s utility in any ambiguous contract that implements action $i\in\{0,\ldots,n+1\}$ is always upper bounded by $R_{i}$.
Consider the classic contract $\langle t,n+1\rangle$, where $t_{0}=0$, and $t_{i}=a_{i}$ for every $i\in[n]$. We claim that this contract is IC and achieves the claimed utility for the principal. Indeed, for every action $i\in[n]$,
$U_{A}({n+1}\mid{t})=\frac{1}{n}\sum_{j=1}^{n}t_{j}-c_{n+1}=\frac{1}{n}A-c_{n+1}=\frac{1}{n}\sum_{j\neq i}t_{j}-(c_{n+1}-\frac{a_{i}}{n})=U_{A}({i}\mid{t}),$
while, for action 00, $U_{A}({n+1}\mid{t})=\frac{1}{n}A-c_{n+1}\geq 0=U_{A}({0}\mid{t})$. Moreover,
$U_{P}({n+1}\mid{t})=\sum_{i=1}^{n}\frac{1}{n}(r_{i}-t_{i})=\sum_{i=1}^{n}\frac{1}{n}(A-a_{i})=A-\sum_{i=1}^{n}\frac{1}{n}a_{i}=\frac{n-1}{n}A,$
as promised. This concludes the proof. ∎
By Proposition 4.2, the optimal $k$-ambiguous contract for $\mathcal{I}(a_{1},\ldots,a_{n})$ implements action $n+1$. Denote this contract by $\langle\tau=\{t^{1},\ldots,t^{k}\},n+1\rangle$. The following proposition demonstrates how to derive an optimal solution to $(n,k)$-Makespan Minimization in instance $\mathcal{J}_{n,k}$ with values $a_{1},\ldots,a_{n}$.
Proposition 4.3.
Given an optimal $k$-ambiguous contract $\langle\tau=\{t^{1},\ldots,t^{k}\},n+1\rangle$ for instance $\mathcal{I}(a_{1},\ldots,a_{n})$, the partition $\mathcal{S}=\{S^{1},\ldots,S^{k}\}$ of $[n]$ given below gives an optimal solution to $(n,k)$-Makespan Minimization in instance $\mathcal{J}_{n,k}$ with values $a_{1},\ldots,a_{n}$. The partition $\mathcal{S}=\{S^{1},\ldots,S^{k}\}$ is given by
$S^{1}=\{i\in[n]\mid U_{A}({i}\mid{t^{1}})\leq U_{A}({n+1}\mid{t^{1}})\}$
$S^{2}=\{i\in[n]\setminus S^{1}\mid U_{A}({i}\mid{t^{2}})\leq U_{A}({n+1}\mid{t^{2}})\}$
$\vdots$
$S^{k}=\left\{i\in[n]\setminus\bigcup_{l\in[k-1]}S^{l}\bigg{|}U_{A}({i}\mid{t^{k}})\leq U_{A}({n+1}\mid{t^{k}})\right\}.$
Before proving Proposition 4.3, we make the following observation.
Observation 4.1.
In instance $\mathcal{I}(a_{1},\ldots,a_{n})$, a payment function $t$ protects action $n+1$ against action $i\in[n]$ if and only if $t_{i}\geq t_{0}+a_{i}$.
Proof.
Contract $t$ protects action $n+1$ against action $i$ if and only if:
$\frac{1}{n}\sum_{j=1}^{n}t_{j}-c_{n+1}\geq\frac{1}{n}\cdot t_{0}+\frac{1}{n}\sum_{j\neq i}t_{j}-(c_{n+1}-\frac{a_{i}}{n}).$
Rearranging, we get $t_{i}\geq t_{0}+a_{i}$, as desired. ∎
We now prove Proposition 4.3.
Proof of Proposition 4.3.
We first show that the expected payment of $\tau$ is at least $\frac{1}{n}\max_{i\in[k]}\sum_{j\in S^{i}}a_{j}$. Since partition $\mathcal{S}$ is constructed such that $t^{i}$ protects action $n+1$ against $S^{i}$, by Proposition 4.1 it holds that for all $i\in[k]$ and $j\in S^{i}$, $t^{i}_{j}\geq t^{i}_{0}+a_{j}\geq a_{j}$, and so the expected payment of $t^{i}$ is at least $\sum_{j\in S^{i}}\frac{1}{n}t^{i}_{j}\geq\frac{1}{n}\sum_{j\in S^{i}}a_{j}.$ Since this is true for all $i\in k$ and $\langle\tau,n+1\rangle$ is consistent, the expected payment of $\tau$ must satisfy
$T_{n+1}(\tau)\geq\frac{1}{n}\max_{i\in[k]}\sum_{j\in S^{i}}a_{j},$ (2)
as desired. Let $\bar{\mathcal{S}}=\{\bar{S}^{1},\ldots,\bar{S}^{k}\}$ be an arbitrary partition of $[n]$, and consider the following set of payment functions, for $i=1,\ldots,k$:
$\bar{t}^{i}_{j}=\begin{cases}0&j=0\text{ or }j\in[n]\setminus\bar{S}^{i},\\ a_{j}&j\in\bar{S}^{i}\\ \end{cases}.$
Note that for all $i\in[k]$, the expected payment of $\bar{t}^{i}$ is $\frac{1}{n}\sum_{j\in\bar{S}^{i}}a_{j}$. By Observation 4.1, it holds that $\bar{t}^{i}$ protects action $n+1$ against $\bar{S}^{i}$. We claim that action $n+1$ is also protected against action 00 by this set of payment functions. Indeed, let $\bar{S}^{i}$ be the set such that $\max_{j\in[n]}a_{j}\in\bar{S}^{i}$. We show that $\bar{t}^{i}$ protects action $n+1$ against action 00. The agent’s utility from action $n+1$ under payment function $\bar{t}^{i}$ is given by
$U_{A}({n+1}\mid{\bar{t}^{i}})=\frac{1}{n}\sum_{j^{\prime}\in\bar{S}^{i}}a_{j^{\prime}}-c_{n+1}=\frac{1}{n}\left(\sum_{j^{\prime}\in\bar{S}^{i}}a_{j^{\prime}}-\max_{j\in[n]}a_{j}\right).$
Since $\sum_{j^{\prime}\in\bar{S}^{i}}a_{j^{\prime}}-\max_{j\in[n]}a_{j}\geq 0$, by definition, and $U_{A}({0}\mid{\bar{t}^{i}})=0$, we conclude that
$U_{A}({n+1}\mid{\bar{t}^{i}})\geq U_{A}({0}\mid{\bar{t}^{i}}),$
as desired. To conclude, the collection of payments $\bar{t}^{1},\ldots,\bar{t}^{k}$ protects action $n+1$ against all other actions.
By Lemma 3.2, we can now apply Algorithm 1 on the set of payment functions $\bar{t}^{1},\ldots,\bar{t}^{k}$ and action $n+1$, to obtain an IC ambiguous contract $\langle\bar{\tau},n+1\rangle$ that implements action $n+1$. The expected payment of $\bar{\tau}$ is the maximum expected payment across payment functions $\bar{t}^{1},\ldots,\bar{t}^{k}$, given by $\frac{1}{n}\max_{i\in[k]}\sum_{j\in\bar{S}^{i}}a_{j}$.
Since $\tau$ is an optimal $k$-ambiguous contract for $\mathcal{I}(a_{1},\ldots,a_{n})$, it holds that $U_{P}({n+1}\mid{\tau})\geq U_{P}({n+1}\mid{\bar{\tau}})$, implying that the expected payment of $\tau$ is at most the expected payment of $\bar{\tau}$. We conclude that
$\frac{1}{n}\max_{i\in[k]}\sum_{j\in S^{i}}a_{j}\leq T_{n+1}(\tau)\leq T_{n+1}(\bar{\tau})=\frac{1}{n}\max_{i\in[k]}\sum_{j\in\bar{S}^{i}}a_{j},$
where the first inequality follows by Equation 2. This concludes the proof. ∎
We are now ready to prove Proposition 4.1.
Proof of Proposition 4.1.
The construction of the instance $\mathcal{I}_{n+2}=\mathcal{I}(a_{1},\ldots,a_{n})$ from the instance $\mathcal{J}_{n,k}$ with values $a_{1},\ldots,a_{n}$ is clearly poly-time. Proposition 4.3 yields a poly-time algorithm for constructing an optimal solution to $\mathcal{J}_{n,k}$ from an optimal $k$-ambiguous contract in $\mathcal{I}_{n+2}$. ∎
5 Ambiguous Contracts of Different Size: Succinctness Gap
In this section, we quantify the loss in principal’s utility due to being restricted to ambiguous contracts with a small number of payment functions. We define the $k$-succinctness gap of an instance as the ratio between the principal’s utility in an optimal $k$-ambiguous contract and an optimal unrestricted ambiguous contract. The $k$-succinctness gap of a family of instances is defined as the worst-case such ratio, over all instances in the family. Formally, let $\mathcal{A}(\mathcal{I}_{n})$ be the set of IC (unrestricted) ambiguous contracts, and $\mathcal{A}_{k}(\mathcal{I}_{n})$ be the set of IC $k$-ambiguous contracts, for an instance $\mathcal{I}_{n}$ with $n$ actions. Then, the $k$-succinctness gap is defined as follows.
Definition 5.1 ($k$-succinctness gap).
Let $\Gamma_{n}$ denote the class of all instances with $n$ actions. The $k$-succinctness gap $\rho_{k}(\mathcal{I}_{n})$ of a given instance $\mathcal{I}_{n}\in\Gamma_{n}$ and the $k$-succinctness gap $\rho_{k}(\Gamma_{n})$ of the class of instances $\Gamma_{n}$, are
$\rho_{k}(\mathcal{I}_{n})=\frac{\max_{\langle\tau,i\rangle\in\mathcal{A}_{k}(\mathcal{I}_{n})}U_{P}(\langle\tau,i\rangle)}{\max_{\langle\tau,i\rangle\in\mathcal{A}(\mathcal{I}_{n})}U_{P}(\langle\tau,i\rangle)}\quad\quad\text{and}\quad\quad\rho_{k}(\Gamma_{n})=\inf_{\mathcal{I}\in\Gamma_{n}}\rho_{k}(\mathcal{I}).$
Note that the succinctness gap $\rho_{k}$ of an instance/a class of instances is less than or equal to one. That is, the closer $\rho_{k}$ is to $1$ the smaller the gap. Moreover, lower bounds on $\rho_{k}$ constitute positive results, while upper bounds on $\rho_{k}$ constitute negative results.
We know that $n-1$ contracts suffice to get the optimal ambiguous contract (see Theorem 2.1). This means that, the $k$-succinctness gap $\rho_{k}(\Gamma_{n})$ for every $k\geq n-1$ is $1$ (i.e., there is no gap). On the other extreme, where the ambiguous contract is composed of a single contract, the succinctness gap coincides with the notion of ambiguity gap studied in Dütting et al. (2024), implying that the succinctness gap $\rho_{1}(\Gamma_{n})$ is exactly $\frac{1}{n-1}$ (see Proposition 4 of their paper). In this section, we establish bounds for intermediate values of $k$.
We first show that, for any $k\in[n-1]$, an ambiguous contract with support size $k$ can obtain at least a fraction $1/(n-k)$ of the optimal ambiguous contract.
Theorem 5.1 (Lower bound, general $k$).
For any $k=1,\ldots,n-1$, the $k$-succinctness gap $\rho_{k}(\Gamma_{n})$ satisfies $\rho_{k}(\Gamma_{n})\geq\frac{1}{n-k}.$
Recall that, for a given action $i$, we use $W_{i}=R_{i}-c_{i}$ to denote the action’s welfare. The proof of Theorem 5.1 uses the following lemma.
Lemma 5.1 (Dütting et al. (2024), Lemma 2).
Consider an instance $\mathcal{I}_{n}\in\Gamma_{n}$ which has a zero-cost action $i$ that yields an expected reward of $R_{i}=0$, then there exists a classic IC contract $\langle t,i^{\prime}\rangle$ such that $U_{P}(\langle t,i^{\prime}\rangle)\geq\frac{1}{n-1}\max_{i\in[n]}W_{i}$.
Proof of Theorem 5.1.
Consider an instance $\mathcal{I}_{n}$. Let $\langle\tau^{\star},a^{\star}\rangle$ be an optimal ambiguous contract for this instance, with principal’s utility $U_{P}^{opt}$. By Theorem 2.1, $\tau^{\star}$ is of size at most $n-1$. Sort all other actions by cost, i.e., $c_{1}\leq c_{2}\leq\ldots\leq c_{n-1}$. Then, by Lemma 5.1, there exists a classic contract $\langle t,a^{\prime}\rangle$ for the instance that includes only actions $a^{\star}$ and $a_{1},\ldots,a_{n-k}$ that achieves principal utility at least $\frac{1}{n-k}W_{a^{\star}}$. The agent’s best response $a^{\prime}$ to payment function $t$ is either action $a^{\star}$ or an action from the set $\{a_{1},\ldots,a_{n-k}\}$. We distinguish between these two cases:
Case 1: The agent’s best response to $t$ is action $a^{\star}$. This means that $t$ protects action $a^{\star}$ against actions $a_{1},\ldots,a_{n-k}$. Moreover, since $\langle\tau^{\star},a^{\star}\rangle$ implements action $a^{\star}$ in the original instance, for all $i\in\{n-k+1,\ldots,n-1$}, there is a payment function $t^{i}\in\tau^{\star}$ that protects action $a^{\star}$ against action $a_{i}$.
The collection of payment functions $\{t,t^{n-k+1},\ldots,t^{n-1}\}$ is a set of at most $1+((n-1)-(n-k+1)+1)=k$ payment functions that contains a corresponding payment function that protects action $a^{\star}$ against each one of actions $a_{1},\ldots,a_{n-1}$. The expected principal’s utility for action $a^{\star}$ under the collection of payment functions $\{t^{n-k+1},\ldots,t^{n-1}\}$ is $U_{P}^{opt}$, and the expected principal’s utility for action $a^{\star}$ under payment function $t$ is at least $\frac{1}{n-k}W_{a^{\star}}\geq\frac{1}{n-k}U_{P}^{opt}$.
Let $\langle\tau,a^{\star}\rangle$ be the ambiguous contract returned by Algorithm 1 for action $a^{\star}$ and the collection of payment functions $\{t,t^{n-k+1},\ldots,t^{n-1}\}$. By Lemma 3.1, this is a IC ambiguous contract with expected principal’s utility of $\min\{U_{P}^{opt},\frac{1}{n-k}W_{a^{\star}}\}\geq\frac{1}{n-k}U_{P}^{opt}$, as required.
Case 2: The agent’s best response to $t$ is action $a_{i}$ for some $i\in[n-k]$. In this case, we know that $t$ protects action $a_{i}$ against $a^{\star}$ and $a_{1},\ldots,a_{n-k}$. Moreover, since $t$ guarantees a principal utility of at least $\frac{1}{n-k}$ of the maximal welfare, and the utility in any contract cannot be higher than the welfare, it holds that $W_{i}\geq U_{P}({i}\mid{t})\geq\frac{1}{n-k}W_{a^{\star}}$.
Consider the payment function $t^{\prime}$ which pays $c_{i}$ for every outcome. This payment function protects action $a_{i}$ against actions $a_{n-k+1},\ldots,a_{n-1}$ because the expected payment under $t^{\prime}$ is $c_{i}$ for all actions, and the costs are sorted so that $c_{i}\leq\min\{c_{n-k+1},\ldots,c_{n-1}\}$. The principal’s utility for action $a_{i}$ under payment function $t^{\prime}$ is $U_{P}({i}\mid{t^{\prime}})=R_{i}-c_{i}=W_{i}$.
Let $\langle\tau,a_{i}\rangle$ be the ambiguous contract returned by Algorithm 1 for action $a_{i}$ and the collection of payment functions $\{t,t^{\prime}\}$. This is an IC ambiguous contract with expected principal’s utility of $\min\{U_{P}({i}\mid{t}),W_{i}\}\geq\frac{1}{n-k}W_{a^{\star}}\geq\frac{1}{n-k}U_{P}^{opt}$, as needed. ∎
Remark 5.1 (Approximation algorithm).
We note that the proof of Theorem 5.1 essentially provides a polynomial-time algorithm that finds an IC $k$-ambiguous contract that gives at least a $\frac{1}{n-k}$ fraction of the optimal principal utility in an unrestricted ambiguous contract. Clearly, this lower bound holds also with respect to the optimal IC $k$-ambiguous contract. Therefore, it also provides a $\frac{1}{n-k}$-approximation for the optimal $k$-ambiguous contract problem.
As a direct corollary of Theorem 5.1, an ambiguous contract of size at most $n-2$ (i.e., just one fewer than the number needed to achieve the optimal utility of an unrestricted ambiguous contract) guarantees at least half of the optimal principal’s utility in the unrestricted case.
The following theorem shows that this is tight. Namely, there exists an instance where no ambiguous contract with strictly less than $n-1$ payment functions yields more than half of the optimal principal’s utility in the unrestricted case.
Theorem 5.2 (Matching upper bound, $k=n-2$).
For $k=n-2$ the $k$-succinctness gap satisfies $\rho_{k}(\Gamma_{n})\leq\frac{1}{2}+O(\frac{1}{n})$.
Proof.
Fix any $n>3$. Let $m=n-1$ and $\epsilon=\frac{1}{n-3}$. Consider the instance given in Figure 3, with $n=m+1$ actions and $m$ outcomes. Note that action $m$ is the zero-cost action with expected reward 00. Note also that
$\frac{m-1}{m-2}=\frac{n-2}{n-3}=1+\frac{1}{n-3}\leq 1+\epsilon.$ (3)
   rewards:    $r_{1}=0$    $r_{2}=0$    $r_{3}=0$    $\dots$    $r_{m}=\frac{(m-1)^{2}}{(m-2)(m-3)}$    costs
action $1$: 00 $\frac{1}{m-1}$ $\frac{1}{m-1}$dots $\frac{1}{m-1}$ $c_{1}=0$
action $2$: $\frac{1}{m-1}$ 00 $\frac{1}{m-1}$dots $\frac{1}{m-1}$ $c_{2}=0$
$\vdots$ $\vdots$ $\vdots$ $\ddots$ $\vdots$ $\vdots$ $\vdots$
action $m$: $\frac{1}{m-1}$ $\frac{1}{m-1}$ $\frac{1}{m-1}$dots 00 $c_{m}=0$
action $m+1$: $\frac{1}{(m-1)^{2}}$ $\frac{1}{(m-1)^{2}}$ $\frac{1}{(m-1)^{2}}$dots $1-\frac{1}{m-1}$ $c_{m+1}=1$
Figure 3: Instance $(c,r,p)$ used in the proof of Theorem 5.2.
We prove the theorem in two parts. We first show that in this instance, there exists an ambiguous contract of size $k=n-1$ that obtains principal utility at least $\frac{2}{m-3}$. We then show that no ambiguous contract of size at most $k=n-2$ can obtain utility greater than $\frac{1+\epsilon}{m-3}$. To prove this, we first show that implementing any action other than action $m+1$ leads to a principal utility of at most $\frac{1+\epsilon}{m-3}$. This is then complemented by showing that no IC ambiguous contract of size at most $k=n-2$ that implements action $m+1$ can obtain principal utility greater than $\frac{1+\epsilon}{m-3}$.
Note that together this shows
$\rho_{n-2}(\Gamma_{n})\leq\frac{\frac{1+\epsilon}{m-3}}{\frac{2}{m-3}}=\frac{1}{2}+\frac{\epsilon}{2}=\frac{1}{2}+O\left(\frac{1}{n}\right),$
as claimed.
Upper bound, $k=n-1$: We start by showing that there exists an ambiguous contract of size $k=n-1=m$ that obtains principal utility at least $\frac{2}{m-3}$.
Consider the collection of payment functions $\tau$ that consists of $n-1$ SOP payment functions $t^{1},\ldots,t^{m}$, where for each $j\in m$, $t^{j}_{j}=\frac{1}{p_{m+1,j}}$ and $t^{j}_{\ell}=0$ for every $\ell\neq j$. For every $j\in[m]$ it holds that: $U_{A}({m+1}\mid{t^{j}})=p_{m+1,j}\cdot\frac{1}{p_{m+1,j}}-c_{m+1}=0$. Therefore, $U_{A}({m+1}\mid{\tau})=0$. Also, for any other action $i\in[m]$ it holds that: $U_{A}({i}\mid{\tau})\leq U_{A}({i}\mid{t^{i}})=0\cdot\frac{1}{p_{m+1,i}}-c_{i}=0$. We have thus shown that $\tau=\{t^{1},\ldots,t^{m}\}$ protects action $m+1$ against all other actions. Note also that, for action $m+1$, the expected payment under $t^{j}$ is $1$ for every $j\in[m]$, so $\langle\tau,m+1\rangle$ is consistent. The argument is completed by noting that the principal’s utility from the ambiguousc contract $\langle\tau,m+1\rangle$ is
$U_{P}(m+1\mid\tau)=R_{m+1}-T_{m+1}(\tau)\geq r_{m}\cdot(1-\frac{1}{m-1})-1=\frac{2}{m-3}.$
Lower bound, $k=n-2$: We next show that no ambiguous contract of size at most $k=n-2=m-1$ can obtain principal utility greater than $\frac{1+\epsilon}{m-3}$.
We first claim that implementing any action $i\neq m+1$ leads to utility of at most $\frac{1+\epsilon}{m-3}$, irrespective of the size of the ambiguous contract. Indeed, for every action $i\neq m+1$ and any ambiguous contract $\langle\tau,i\rangle$, it holds that
$U_{P}({i}\mid{\tau})\leq R_{i}=\frac{1}{m-1}\sum_{j=1}^{m-1}j+\frac{r_{m}}{m-1}=\frac{m-1}{(m-2)(m-3)}\leq\frac{1+\epsilon}{m-3},$
where the first inequality follows since the principal’s utility from implementing any action is upper bounded by the action’s welfare, and the last inequality follows by Equation (3).
We next show that the principal’s utility from any IC ambiguous contract $\langle\tau,m+1\rangle$ of size at most $k=n-2=m-1$ that implements action $m+1$ is at most $\frac{1+\epsilon}{m-3}$.
Let $\theta$ denote the expected payment for action $m+1$ under $\tau$, and let $t$ be an arbitrary payment function in $\tau$. Since the expected payment for action $m+1$ under $\tau$ is $\theta$, there must exist $j_{t}\in[m]$ for which $t_{j_{t}}\geq\theta$. For every action $i\in[m]\setminus\{j_{t}\}$ it holds that $p_{i,j_{t}}=\frac{1}{m-1}$, thus the expected payment for action $i$ under $t$ is at least $t_{j_{t}}\cdot\frac{1}{m-1}\geq\frac{\theta}{m-1}$. Combined with the fact that $\tau$ is of size at most $m-2$, it follows that |$|\{j_{t}\mid t\in\tau\}|\leq m-2$. Therefore, there must exist an action $i\in[n]$ that is not in $\{j_{t}\mid t\in\tau\}$. The expected payment for that action $i$ under every $t\in\tau$ is at least $\frac{\theta}{m-1}$.
It follows that, in order for $\tau$ to protect action $m+1$ against action $i$, it must hold that
$\theta-c_{m+1}=U_{A}({m+1}\mid{\tau})\geq U_{A}({i}\mid{\tau})\geq\frac{\theta}{m-1}.$
Substituting $c_{m+1}=1$ and rearranging, we get that $\theta\geq\frac{m-1}{m-2}$. It follows that
$U_{P}({m+1}\mid{\tau})=R_{m+1}-\theta\leq\frac{m-1}{m-3}-\frac{m-1}{m-2}=\frac{m-1}{(m-3)(m-2)}\leq\frac{1+\epsilon}{m-3},$
where the last inequality follows from Equation (3). ∎
We next establish an upper bound of $\frac{k}{n+k-2}$ on the succinctness gap for ambiguous contracts with support of size at most $k$ for $k=1,\ldots,n-2$. The proof is given in Appendix C.
Theorem 5.3 (Upper bound, general $k$).
For every integer $n$ of the form $k(\tilde{n}-2)+2$, for some integer $\tilde{n}\geq 3$, there exists an instance $\mathcal{I}_{n}$, such that $\rho_{k}(\mathcal{I}_{n})\leq\frac{k}{n+k-2}$.
Notably, together with the lower bound in Theorem 5.1, this generalizes the result in Dütting et al. (2024a), which shows that the ambiguity gap, defined as the worst-case ratio between unrestricted ambiguous contracts and classic contracts, is $n-1$. This is because for $k=1$ we get an upper bound of $\frac{k}{n+k-2}=\frac{1}{n-1}$, and a matching lower bound of $\frac{1}{n-k}=\frac{1}{n-1}$. In addition, for $k=n-2$ the bound is $\frac{n-2}{n+n-2-2}=\frac{1}{2}$, matching the bound in Theorem 5.2.
6 Conclusion and Future Work
[…]
Our work raises several interesting directions for future work. One question is how the structure of optimal succinct ambiguous contracts changes when a monotonicity constraint is imposed. On the computational side, exploring polynomial-time approximation algorithms is another promising direction.
Appendix A Proofs Omitted from Section 3
Proof of Lemma 3.1.
We begin by showing that $\langle\tau,i\rangle$ is incentive compatible. We first show that $\langle\tau,i\rangle$ is consistent. To this end, observe that the expected payment for action $i$ under any payment function $\bar{t}^{j}\in\tau$ is
$\bar{T}_{i}^{j}=\sum_{\ell\in[m]}p_{i\ell}\bar{t}^{j}_{\ell}=\sum_{\ell\in[m]}p_{i\ell}(t^{j}+(\theta-T^{j}_{i}))_{\ell}=\theta-T^{j}_{i}+\sum_{\ell\in[m]}p_{i\ell}t^{j}_{\ell}=\theta.$ (4)
Next we show that action $i$ is a best response to $\tau$. Fix an action $i^{\prime}\neq i$, and let $j$ be the index such that $i^{\prime}\in S^{j}$ We have
$U_{A}({i^{\prime}}\mid{\bar{t}^{j}})$ $=\sum_{\ell\in[m]}p_{i^{\prime}\ell}\bar{t}^{j}_{\ell}-c_{i^{\prime}}=\sum_{\ell\in[m]}p_{i^{\prime}\ell}(t^{j}_{\ell}+(\theta-T^{j}_{i}))-c_{i^{\prime}}=T^{j}_{i^{\prime}}-c_{i^{\prime}}+\theta-T^{j}_{i},\quad\text{and}$ (5)
$U_{A}({i}\mid{\bar{t}^{j}})$ $=\bar{T}^{j}_{i}-c_{i}=\theta-c_{i},$ (6)
where the last equality holds by (4).
On the other hand, since $t^{j}$ protects action $i$ against the set of actions $S^{j}$, we have that
$T^{j}_{i}-c_{i}=U_{A}({i}\mid{t^{j}})\geq U_{A}({i^{\prime}}\mid{t^{j}})=T^{j}_{i^{\prime}}-c_{i^{\prime}}.$ (7)
It follows that
$U_{A}({i}\mid{\tau})-U_{A}({i^{\prime}}\mid{\tau})\geq U_{A}({i}\mid{\bar{t}^{j}})-U_{A}({i^{\prime}}\mid{\bar{t}^{j}})=(T^{j}_{i}-c_{i})-(T^{j}_{i^{\prime}}-c_{i^{\prime}})\geq 0.$
For the first inequality, we use that $U_{A}({i}\mid{\tau})=U_{A}({i}\mid{\bar{t}^{j}})$, by the consistency of $\tau$, and that $U_{A}({i^{\prime}}\mid{\tau})\leq U_{A}({i^{\prime}}\mid{\bar{t}^{j}})$ since $\bar{t}^{j}\in\tau$. The equality holds by equations (5) and (6). The final inequality holds by (7).
For the claim regarding the principal’s utility note that for any $\bar{t}^{j}\in\tau$ it holds that
$U_{P}({i}\mid{\bar{t}^{j}})=R_{i}-\theta=R_{i}-\max\{T_{i}^{1},\ldots,T_{i}^{k}\}=\min_{j\in[k]}\{R_{i}-T_{i}^{j}\}=\min_{j\in[k]}\{U_{P}({i}\mid{t^{j}})\},$
where the first equality follows by (4). ∎
Proof of Lemma 3.2.
Consider first the case where Algorithm 2 returns an ambiguous contract $\langle\tau,i\rangle$. By construction, the collection of payment functions $\{t^{1},\ldots,t^{k}\}$ determined by the algorithm in the for-loop protects action $i$ against partition $\mathcal{S}$. So, by Lemma 3.1, the ambiguous contract $\langle\tau,i\rangle$ returned by Algorithm 2 is a (consistent) $k$-ambiguous contract that implements action $i$. Moreover, with $\theta=\max_{j\in[k]}T_{i}^{j}$ defined as in Algorithm 1, we have $\tau=\{\bar{t}^{1},\ldots,\bar{t}^{k}\}$, where for each $\bar{t}^{j}$ it holds that $\bar{t}^{j}=t^{j}+(\theta-T_{i}^{j})\cdot\vec{1}=t^{j}+c^{j}\cdot\vec{1}$. So each $\bar{t}^{j}\in\tau$ corresponds to a shifted min-pay contract $\langle\bar{t}^{j},i\rangle$ for the subinstance $\mathcal{I}_{\{i\}\cup S^{j}}$ obtained by restricting the original instance to actions $\{i\}\cup S^{j}$.
It remains to show that $\langle\tau,i\rangle$ is an optimal IC $k$-ambiguous contract with respect to partition $\mathcal{S}$. To this end, consider an arbitrary IC $k$-ambiguous contract $\langle\hat{\tau}=\{\hat{t}^{1},\ldots,\hat{t}^{k}\},i\rangle$ such that $\hat{\tau}=\{\hat{t}^{1},\ldots,\hat{t}^{k}\}$ protects action $i$ against partition $\mathcal{S}$. We show that $U_{P}({i}\mid{\hat{\tau}})\leq U_{P}({i}\mid{\tau})$. Let $j^{\star}\in\operatorname*{arg\,max}_{j\in[k]}T_{i}^{j}$. For this $j^{\star}$, it holds that $\theta=T_{i}^{j^{\star}}$, and thus $\bar{t}^{j^{\star}}=t^{j^{\star}}+(\theta-T^{j^{\star}}_{i})\cdot\vec{1}=t^{j^{\star}}$. It follows that $\bar{t}^{j^{\star}}=t^{j^{\star}}$ is an optimal solution to MIN-PAY-LP$(i)$ on the subinstance $\mathcal{I}_{\{i\}\cup S^{j^{\star}}}$. Since $\hat{t}^{j^{\star}}$ is also a feasible solution to MIN-PAY-LP$(i)$ on the subinstance $\mathcal{I}_{\{i\}\cup S^{j^{\star}}}$, it holds that $U_{P}({i}\mid{\hat{t}^{j^{\star}}})\leq U_{P}({i}\mid{t^{j^{\star}}}).$ Combined with the fact that both $\tau$ and $\hat{\tau}$ are consistent, we get:
$U_{P}({i}\mid{\hat{\tau}})=U_{P}({i}\mid{\hat{t}^{j^{\star}}})\leq U_{P}({i}\mid{t^{j^{\star}}})=U_{P}({i}\mid{\tau}).$
Consider next the case where Algorithm 2 returns Null. The algorithm returns Null only if there exists $j\in[k]$ for which running MIN-PAY-LP$(i)$ on the subinstance $\mathcal{I}_{\{i\}\cup S^{j}}$ returns Null. But in this case, there is no payment function that protects action $i$ against $S^{j}$, concluding the argument. ∎
Proof of Proposition 3.1.
$\Rightarrow$: Assume action $i$ is implementable by a $k$-ambiguous contract, and let $\langle\tau=\{t^{1},\ldots,t^{k}\},i\rangle$ be a contract of size $k$ implementing action $i$. Define the following sets:
$S^{1}=\{i^{\prime}\in[n]\setminus\{i\}\mid U_{A}({i^{\prime}}\mid{t^{1}})\leq U_{A}({i}\mid{t^{1}})\}$
$S^{2}=\{i^{\prime}\in\left([n]\setminus\{i\}\right)\setminus S^{1}\mid U_{A}({i^{\prime}}\mid{t^{2}})\leq U_{A}({i}\mid{t^{2}})\}$
$\vdots$
$S^{k}=\left\{i^{\prime}\in\left([n]\setminus\{i\}\right)\setminus\bigcup_{l\in[k-1]}S^{l}\bigg{|}U_{A}({i^{\prime}}\mid{t^{k}})\leq U_{A}({i}\mid{t^{k}})\right\}$
and let $\mathcal{S}=\{S^{1},S^{2},\ldots,S^{k}\}$. By design, for all $j\in[k]$, it holds that $t^{j}$ protects action $i$ against $S^{j}$. Therefore, $t^{j}$ is a solution, not necessarily optimal, to MIN-PAY-LP$(i)$ in the subinstance $\mathcal{I}_{\{i\}\cup S^{j}}$, and in particular, this LP is feasible. From Proposition 2.1, it holds that there does not exist a convex combination $\lambda_{i^{\prime}}\in[0,1]$ of the actions $i^{\prime}\in S^{j}$ that yields the same distribution over rewards $\sum_{i^{\prime}\in S^{j}}\lambda_{i^{\prime}}p_{i^{\prime}j}=p_{ij}$ for all $j$ but at a strictly lower cost $\sum_{i^{\prime}\in S^{j}}\lambda_{i^{\prime}}c_{i^{\prime}}<c_{i}$.
$\Leftarrow$: Assume there exists a partition $\mathcal{S}$ of $[n]\setminus\{i\}$ to $k$ sets, such that for every set $S\in\mathcal{S}$, action $i$ is implementable by a classic contract in $\mathcal{I}_{S\cup\{i\}}$. In this case, let $\{\hat{t}^{1},\ldots,\hat{t}^{k}\}$ be a collection of payment function corresponding to solutions of MIN-PAY-LP$(i)$ in the subinstances $\mathcal{I}_{\{i\}\cup S^{1}},\ldots,\mathcal{I}_{\{i\}\cup S^{k}}$. By running Algorithm 1 with the input $(i,\{\hat{t}^{1},\ldots,\hat{t}^{k}\})$, we obtain an IC ambiguous contract $\langle\hat{\tau}=\{\hat{t}^{1},\ldots,\hat{t}^{k}\},i\rangle$ of size $k$ implementing action $i$, as needed. ∎
Appendix B NP-Hardness of  Makespan Minimization
In this appendix, we prove the following theorem, which establishes NP-hardness of the $(n,k)$-Makespan Minimization problem. The proof is by reduction from the partition problem (e.g., Garey and Johnson, 1979, Section 3.1.5). An instance of this problem is given by a set of $n$ positive integers $a_{1},\ldots a_{n}$. The question is whether there exists a partition of the set into two subsets with equal sum, namely, whether there exists a set $S\subseteq[n]$ such that
$\sum_{i\in S}a_{i}=\sum_{i\notin S}a_{i}.$
Theorem B.1.
The $(n,k)$-Makespan Minimization problem is NP-hard for any $n,k$ such that $n=\textsf{poly}(n-k)$. This includes the case where $k=c$ for some constant $c$, and the case where $k=\beta\cdot n$ for some $\beta\in(0,1).$
Proof of Theorem B.1.
The proof is based on a reduction from partition to $(n,k)$-Makespan Minimization. Let $b_{1},\ldots,b_{m}$ be an instance of the partition problem with size $m$, and denote the sum by $2B=\sum_{i=1}^{m}b_{i}$. For $k\geq 2$ we define the following instance of $(n,k)$-Makespan Minimization: Let $n=m+k-2$ and define the multiset $A=\{a_{1},\ldots,a_{n}$} such that:
$a_{i}=\begin{cases}b_{i}&i\in[m]\\ B&i>m\\ \end{cases}.$
We claim that the partition instance is a YES instance if and only if the $(n,k)$-Makespan Minimization instance has a solution such that the minimum value is $B$ (note that there cannot be a minimum value lower than $B$ since there are elements with this size).
Assume that the partition instance is a YES instance, and denote by $S$ the set of indices such that $\sum_{i\in S}b_{i}=\sum_{i\in[m]\setminus S}b_{i}=B$. Consider the following partition of $A$ for the Makespan Minimization: $\mathcal{A}=\{A_{1},\ldots,A_{k}\}$ where $A_{1}=\{a_{i}\mid i\in S\}$, $A_{2}=\{a_{i}\mid i\in[m]\setminus S\}$ and for $j=3,\ldots,k$, $A_{j}=\{a_{j}\}$. It holds that:
$\sum_{i\in A_{1}}a_{i}=\sum_{i\in S}b_{i}=B,$
$\sum_{i\in A_{2}}a_{i}=\sum_{i\in[m]\setminus S}b_{i}=B,\qquad\text{and}$
$\sum_{i\in A_{3}}a_{i}=\ldots=\sum_{i\in A_{k}}a_{i}=B.$
Hence the makespan is $\max_{j\in[k]}\sum_{a\in A_{j}}a=B$.
Next, assume there exists a partition of $A$, denoted by $\mathcal{A}=\{A_{1},\ldots,A_{k}\}$, such that the makespan is $\max_{j\in[k]}\sum_{a\in A_{j}}a=B$. Since we have $k-2$ elements of size $B$ and all other elements are positive, all elements of size $B$ need to be in a separate set in the partition $\mathcal{A}$ of $A$. This leaves two sets and the first $m$ elements of $A$, which, by construction, coincide with $b_{1},\ldots,b_{m}$ and sum to $2B$. Therefore the last two sets in $\mathcal{A}$ must part the first $m$ elements of $A$, which equal the elements of the instance of partition, to two sets that sum to $B$.
In order for the reduction to be polynomial in $m$, we need $n=m+k-2=\textsf{poly}(m)$, so $n,k$ must satisfy $n=\textsf{poly}(n-k)$. This is satisfied when $k=c$ for some constant $c$ since then $n=\textsf{poly}(n-k)=\textsf{poly}(n-c)$. It also holds for any $k=\beta\cdot n$ for some $\beta\in(0,1)$ since then $n=\textsf{poly}(n-k)=\textsf{poly}((1-\beta)n)$. ∎
Appendix C Proofs Omitted from Section 5
Proof of Theorem 5.3.
Fix an integer $\tilde{n}\geq 3$. Let $n=k(\tilde{n}-2)+2$, let $\gamma,\epsilon\in(0,1)$, and let $\delta=\frac{1}{k}\epsilon\cdot\gamma^{\tilde{n}-2}$. Consider the instance given in Figure 4 with $n$ actions and $k+2$ outcomes.
Action $1$ is a zero-cost and zero-reward action. Then there are $k$ groups of actions $A_{j}=\{i_{j}\mid 2\leq i\leq\tilde{n}-1\}$ such that action $i_{j}$ has a cost of $c_{i_{j}}=\frac{1}{\gamma^{i-2}}-(i-1)+(i-2)\gamma$ (note that the cost only depends on $i$) and puts a probability of $2\delta$ on outcomes in $[k]\setminus\{j\}$, a probability of $1-\gamma^{\tilde{n}-i}-(2k-2)\delta$ on outcome $k+1$, and a probability of $\gamma^{\tilde{n}-i}$ on outcome $k+2$. The final action $\tilde{n}$ will be our target action. That action has a cost of $c_{\tilde{n}}=\frac{1}{\gamma^{\tilde{n}-2}}-(\tilde{n}-1)+(\tilde{n}-2)\gamma$ and puts a probability $\delta$ on outcomes in $[k]$ and a probability of $1-k\delta$ on outcome $k+2$.
rewards: $r_{1}$ $r_{2}$ldots $r_{k}$ $r_{k+1}$ $r_{k+2}$
00 00 00 00 $\frac{1}{\gamma^{\tilde{n}-2}}$ costs
action $1$: $\frac{1}{k}$ $\frac{1}{k}$ldots $\frac{1}{k}$ 00 00 $c_{1}=0$
action $i_{1}$: 00 $2\delta$ldots $2\delta$ $1-\gamma^{\tilde{n}-i}-(2k-2)\delta$ $\gamma^{\tilde{n}-i}$ $c_{i_{1}}=\frac{1}{\gamma^{i-2}}-(i-1)+(i-2)\gamma$
$12\leq i\leq\tilde{n}-$ 1
action $i_{2}$: $2\delta$ 00 $\ldots$ $2\delta$ $1-\gamma^{\tilde{n}-i}-(2k-2)\delta$ $\gamma^{\tilde{n}-i}$ $c_{i_{2}}=\frac{1}{\gamma^{i-2}}-(i-1)+(i-2)\gamma$
$12\leq i\leq\tilde{n}-$ 1
$\vdots$ $\vdots$ $\vdots$ $\ddots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$
action $i_{k}$: $2\delta$ $2\delta$ldots 00 $1-\gamma^{\tilde{n}-i}-(2k-2)\delta$ $\gamma^{\tilde{n}-i}$ $c_{i_{k}}=\frac{1}{\gamma^{i-2}}-(i-1)+(i-2)\gamma$
$12\leq i\leq\tilde{n}-$ 1
action $\tilde{n}$: $\delta$ $\delta$ldots $\delta$ 00 $1-k\delta$ $c_{\tilde{n}}=\frac{1}{\gamma^{\tilde{n}-2}}-(\tilde{n}-1)+(\tilde{n}-2)\gamma$
Figure 4: Instance $(c,r,p)$ used in the proof of Theorem 5.3.
We will first show that with an ambiguous contract of size $k+1$, the principal can achieve a utility of $\tilde{n}-1$ by implementing action $\tilde{n}$. Then we will show that the principal’s utility from any ambiguous contract of size at most $k$ that implements action $\tilde{n}$ is at most $1$. For this we will show that the principal’s utility for action $\tilde{n}$ under any payment function $t$ that protects against any pair of actions from $\{(\tilde{n}-1)_{1},(\tilde{n}-1)_{2},\ldots,(\tilde{n}-1)_{k},1\}$ is at most $1$. Finally, we will show that the principal’s utility under any ambiguous contract that implements an action other than $\tilde{n}$ is at most $1$. The bound on the succinctness gap follows since
$\frac{1}{\tilde{n}-1}=\frac{k}{n+k-2}.$
An ambiguous contract of size $k+1$ can achieve a principal’s utility of $\tilde{n}-1$. We show that action $\tilde{n}$ can be implemented with an ambiguous contract of size $k+1$ with an expected payment equal to $c_{\tilde{n}}$. The ambiguous contract that achieves this is $\langle\{t^{1},t^{2},\ldots,t^{k},t^{k+1}\},\tilde{n}\rangle$ with
$t^{1}=(\frac{c_{\tilde{n}}}{\delta},0,\ldots,0,0,0),$
$t^{2}=(0,\frac{c_{\tilde{n}}}{\delta},\ldots,0,0,0),$
$\ldots$,… ,
$t^{k}=(0,0,\ldots,\frac{c_{\tilde{n}}}{\delta},0,0),\qquad\quad\text{and}$
$t^{k+1}=(0,0,\ldots,0,0,\frac{c_{\tilde{n}}}{1-k\delta}).$
It is easy to verify that this contract is consistent, and entails an expected payment of $c_{\tilde{n}}$ for action $\tilde{n}$. It is IC, because for all actions $i\neq\tilde{n}$ it gives a minimum payment of zero.
The claim about the principal’s utility follows, as the welfare from action $\tilde{n}$ is $W_{\tilde{n}}=(\tilde{n}-1)-(\tilde{n}-2)\gamma-\epsilon$, and $W_{\tilde{n}}\rightarrow\tilde{n}-1$ as $\gamma,\epsilon\rightarrow 0$.
A payment function $t$ protecting action $\tilde{n}$ from action $1$ and action $(\tilde{n}-1)_{i}$. We show that if a payment function $t$ protects action $\tilde{n}$ against both action $1$ and action $(\tilde{n}-1)_{1}$, then $U_{P}({\tilde{n}}\mid{t})\leq 1$. An analogous argument applies to any payment function $t$ that protects action $\tilde{n}$ against action $1$ and action $(\tilde{n}-1)_{i}$ with $i\neq 1$, and shows $U_{P}({\tilde{n}}\mid{t})\leq 1$.
Consider action $\tilde{n}$, and any payment function $t$ that protects action $\tilde{n}$ against both action $1$ and action $(\tilde{n}-1)_{1}$. Since action $\tilde{n}$ places zero probability on outcome $k+1$, we can without loss of generality assume $t_{k+1}=0$. To obtain an upper bound on the principal’s utility, we compare action $\tilde{n}$ to action $1$ and action $(\tilde{n}-1)_{1}$. Since $t$ protects action $\tilde{n}$ against action $1$, it follows that
$\delta(t_{1}+t_{2}+\ldots+t_{k})+(1-k\delta)t_{k+2}-c_{\tilde{n}}\geq\frac{1}{k}(t_{1}+t_{2}+\ldots+t_{k}),$
or equivalently
$t_{k+2}-\frac{1}{k}(t_{1}+t_{2}+\ldots+t_{k})\geq\frac{c_{\tilde{n}}}{1-k\delta}.$
An important consequence of this is that
$t_{k+2}-\frac{1}{k}(t_{1}+t_{2}+\ldots+t_{k})\geq 0.$ (8)
On the other hand, since the agent does not want to deviate to action $(\tilde{n}-1)_{1}$, it must hold that
$\delta(t_{1}+t_{2}+\ldots+t_{k})+(1-k\delta)t_{k+2}-c_{\tilde{n}}\geq\gamma t_{k+2}+2\delta(t_{2}+\ldots+t_{k})-c_{(\tilde{n}-1)_{1}}\geq\gamma t_{k+2}-c_{(\tilde{n}-1)_{1}}.$
It must therefore hold that
$\gamma t_{k+2}-c_{(\tilde{n}-1)_{1}}$ $\leq\delta(t_{1}+t_{2}+\ldots+t_{k})+(1-k\delta)t_{k+2}-c_{\tilde{n}}$
$=t_{k+2}-c_{\tilde{n}}-k\delta(t_{k+2}-\frac{1}{k}(t_{1}+t_{2}+\ldots+t_{k}))\leq t_{k+2}-c_{\tilde{n}},$
where the second inequality holds by Equation (8). Rearranging and substituting $c_{\tilde{n}}$ and $c_{(\tilde{n}-1)_{1}}$, this yields
$t_{k+2}\geq\frac{1}{1-\gamma}(c_{\tilde{n}}-c_{(\tilde{n}-1)_{1}})=\frac{1}{\gamma^{\tilde{n}-2}}-1.$
Hence, the principal’s utility from action $\tilde{n}$ under payment function $t$ is at most
$(1-k\delta)r_{k+2}-(1-k\delta)t_{k+2}\leq 1-k\delta,$
as claimed.
A payment function $t$ protecting action $\tilde{n}$ from action $(\tilde{n}-1)_{i}$ and action $(\tilde{n}-1)_{j}$. We next show that if a payment function $t$ protects action $\tilde{n}$ against both action $(\tilde{n}-1)_{1}$ and action $(\tilde{n}-1)_{2}$, then $U_{P}({\tilde{n}}\mid{t})\leq 1$. An analogous argument applies to any payment function $t$ that protects action $\tilde{n}$ against two actions of the form $(\tilde{n}-1)_{i}$ and $(\tilde{n}-1)_{j}$ for $i\neq j$, and shows $U_{P}({\tilde{n}}\mid{t})\leq 1$.
Consider action $\tilde{n}$, and any payment function $t$ that protects action $\tilde{n}$ against both action $(\tilde{n}-1)_{1}$ and action $(\tilde{n}-1)_{2}$. As before, we can assume without loss of generality that $t_{k+1}=0$. Since $t$ protects action $\tilde{n}$ against action $(\tilde{n}-1)_{1}$, it holds that
$\delta(t_{1}+t_{2}+\ldots+t_{k})+(1-k\delta)t_{k+2}-c_{\tilde{n}}\geq\gamma t_{k+2}+2\delta(t_{2}+\ldots+t_{k})-c_{(\tilde{n}-1)_{1}}.$
It must therefore hold that
$\delta(t_{1}+t_{2})+(1-k\delta)t_{k+2}-c_{\tilde{n}}$ $\geq\gamma t_{k+2}+2\delta t_{2}+\delta(t_{3}+\ldots+t_{k})-c_{(\tilde{n}-1)_{1}}$
$\geq\gamma t_{k+2}+2\delta t_{2}-c_{(\tilde{n}-1)_{1}},$
where we used that $\delta\geq 0$ and $t_{j}\geq 0$ for all $j$. Similarly, since $t$ protects action $\tilde{n}$ against action $(\tilde{n}-1)_{2}$, it holds that
$\delta(t_{1}+t_{2})+(1-k\delta)t_{k+2}-c_{\tilde{n}}\geq\gamma t_{k+2}+2\delta t_{1}-c_{(\tilde{n}-1)_{2}}.$
By adding both inequalities, we obtain
$2\delta(t_{1}+t_{2})+2(1-k\delta)t_{k+2}-2c_{\tilde{n}}\geq 2\gamma t_{k+2}+2\delta(t_{1}+t_{2})-2c_{(\tilde{n}-1)_{1}},$
where we used that $c_{(\tilde{n}-1)_{1}}=c_{(\tilde{n}-1)_{2}}$.
Rearranging this inequality we obtain
$\gamma t_{k+2}-c_{(\tilde{n}-1)_{1}}\leq(1-k\delta)t_{k+2}-c_{\tilde{n}}\leq t_{k+2}-c_{\tilde{n}}.$
Solving for $t_{k+2}$ and substituting $c_{\tilde{n}}$ and $c_{(\tilde{n}-1)_{1}}$, we obtain
$t_{k+2}\geq\frac{1}{1-\gamma}(c_{\tilde{n}}-c_{(\tilde{n}-1)_{1}})=\frac{1}{\gamma^{\tilde{n}-2}}-1.$
Hence the principal’s utility from action $\tilde{n}$ is at most
$(1-k\delta)r_{k+2}-(1-k\delta)t_{k+2}\leq 1-k\delta,$
as needed.
An ambiguous contract of size at most $k$ implementing action $\tilde{n}$. Combining the two previous claims, we conclude that any ambiguous contract of size at most $k$, implementing action $\tilde{n}$ can achieve a principal’s utility of at most $1$. This is because in an IC ambiguous contract $\langle\tau,\tilde{n}\rangle$ of size at most $k$, there must be a payment function $t\in\tau$ that protects action $\tilde{n}$ against at least a single pair from $\{(\tilde{n}-1)_{1},(\tilde{n}-1)_{2},\ldots,(\tilde{n}-1)_{k},1\}$. Therefore,
$U_{P}({\tilde{n}}\mid{\tau})=U_{P}({\tilde{n}}\mid{t})\leq 1,$
where the equality holds by consistency.
An ambiguous contract implementing any action other than action $\tilde{n}$. We next show that for any ambiguous contract (of any size), that implements any action other than action $\tilde{n}$, the principal’s utility is upper bounded by $1$. To this end, consider an IC ambiguous contract $\langle\tau,i^{\star}\rangle$ with $i^{\star}\neq\tilde{n}$. We distinguish between two cases.
Case 1: Contract $\langle\tau,i^{\star}\rangle$ implements action $1$, or an action of the form $2_{j}$. These actions have welfare of at most $1$, implying that the principal’s utility cannot exceed $1$.
Case 2: Contract $\langle\tau,i^{\star}\rangle$ implements an action of type $i_{j}$ for some $i\geq 3$, $j\in[k]$. For simplicity, we will show it for $i_{1}$. Suppose $t\in\tau$ protects action $i_{1}$ against action $(i-1)_{1}$. Since action $i_{1}$ places zero probability on outcome $1$, we can assume without loss of generality that $t_{1}=0$.
According to this constraint, we must have
$2\delta(t_{2}+\ldots+t_{k})+$ $(1-\gamma^{\tilde{n}-i}-2(k-1)\delta)t_{k+1}+\gamma^{\tilde{n}-i}t_{k+2}-c_{i_{1}}\geq$
$2\delta(t_{2}+\ldots+t_{k})+(1-\gamma^{\tilde{n}-(i-1)}-2(k-1)\delta)t_{k+1}+\gamma^{\tilde{n}-(i-1)}t_{k+2}-c_{(i-1)_{1}},$
or, equivalently,
$(1-\gamma^{\tilde{n}-i}-2(k-1)\delta)t_{k+1}+\gamma^{\tilde{n}-i}t_{k+2}-c_{i_{1}}\geq(1-\gamma^{\tilde{n}-i+1}-2(k-1)\delta)t_{k+1}+\gamma^{\tilde{n}-i+1}t_{k+2}-c_{(i-1)_{1}}.$
Since $1-\gamma^{\tilde{n}-i}-2(k-1)\delta<1-\gamma^{\tilde{n}-i+1}-2(k-1)\delta$ and $\gamma^{\tilde{n}-i}>\gamma^{\tilde{n}-i+1}$, we obtain a lower bound on the expected payment $T_{i_{1}}(t)$ by setting $t_{k+1}=0$ and finding the smallest $t_{k+2}$ such that
$\gamma^{\tilde{n}-i}t_{k+2}-c_{i_{1}}\geq\gamma^{\tilde{n}-i+1}t_{k+2}-c_{(i-1)_{1}}.$
Rearranging and substituting $c_{i_{1}}$ and $c_{(i-1)_{1}}$, this yields
$t_{k+2}\geq\frac{1}{\gamma^{\tilde{n}-i}}\cdot\frac{c_{i_{1}}-c_{(i-1)_{1}}}{1-\gamma}=\frac{1}{\gamma^{\tilde{n}-i}}\left(\frac{1}{\gamma^{i-2}}-1\right).$
Thus, the principal’s utility from action $i_{1}$ is at most
$\gamma^{\tilde{n}-i}r_{k+2}-\gamma^{\tilde{n}-i}t_{k+2}\leq\frac{1}{\gamma^{i-2}}-\left(\frac{1}{\gamma^{i-2}}-1\right)=1,$
which concludes the proof.